### v2

# A Cross-Cultural Analysis of Social Norms in Bollywood and Hollywood Movies

Abstract
Understanding how social norms vary across cultures can help us build culturally aligned NLP systems. We propose a culture-agnostic approach to norm discovery, using moral emotions, shame and pride, to identify examples of normative expectations and extract corresponding social norms. We present the first cross-cultural self-conscious emotions dataset, obtained from  5.4K Bollywood and Hollywood movies, along with over 10K extracted social norms. We validate our dataset using native speakers and demonstrate how our dataset reveals variations in social norms that align with the cultural dichotomy observed in these nations -- e.g., Bollywood movies emphasize shame due to deviation from social roles, and express pride in family honor, while Hollywood shames poverty and incompetence, and takes pride in ethical behavior. Notably, females are shamed more across both cultures and both cultures shame women for violating similar normative expectations.
A Cross-Cultural Analysis of Social Norms in Bollywood and Hollywood Movies
1 Introduction
Normative expectations, [^1] Bicchieri (2005) uses the term normative expectations, Lahti and Weinstein (2005) use moral norms, and many social psychologists Cialdini et al. (1990) use injunctive norms. We use the terms social norms and normative expectations interchangeably. commonly called ‘‘social norms’’, refer to people’s beliefs about what others approve of or expect people to do. These norms play a significant role in shaping and regulating social behavior and can reveal underlying moral and ethical values.
Recent research on social alignment of LLMs has either crowd-sourced social norms from QnA style posts on social media platforms or prompted pre-trained LLMs to extract them Jiang et al. (2021); Nahian et al. (2020); Fung et al. (2022). However, a majority of these studies are conducted in English and capture Western social situations. Furthermore, LLMs such as GPT-4 OpenAI (2023) are predominantly trained on data from Western cultures, Havaldar et al. (2023b), and thus, the generated "social norms" tend to be Anglocentric.
Look, Alok is in love with me and I love him too.
Shameless one! Have you lost all your shame?
What is there to be ashamed in this? Don’t sisters love their brothers in this vicinity?
Okay, I am going.
$\Rightarrow$ Incestuous relationship evokes shame
Sister-in-law! Congrats, sister-in-law! Big brother has started working!
Really?
Now you will have a place of pride in this family!
Yes, please!
Transfer 10 lakhs, okay? Bye!
$\Rightarrow$ Employed husband evokes pride
Table 1: Excerpts of dialogues with target search words.
Figure 1: An overview of our approach comprising two key steps (a) Vocabulary approach and (b) Prompting a pre-trained LLM.
Self-conscious emotions, or moral emotions, are emotional reactions to a violation or an observance of normative expectations in society. In turn, these emotions reflect social sanctions and approvals. Unlike basic emotions such as joy and anger, self-conscious emotions such as shame and pride are evoked by self-reflection and self-evaluation. Examining situations evoking self-conscious emotions can thus reveal unspoken social norms specific to a culture (See Table 1). While the function (i.e., social regulation) of self-conscious emotions is consistent across cultures Goetz and Keltner (2007); Fessler (2007); Schaumberg and Skowronek (2022), their appraisals and expressions vary Wong and Tsai (2007); Lewis et al. (2010); Wegge et al. (2022).
In this paper, we (a) examine cross-cultural variations in linguistic expressions of shame and pride, (b) propose a culture-agnostic approach for norm discovery and release a dataset of $10K$ normative expectations and, (c) examine cross-cultural variations in social norms. We study India and the United States of America to investigate the utility of moral emotions for cross-culture norm discovery. India is predominantly a collectivist society where one’s sense of self is interwoven with community beliefs, whereas the USA is predominantly an individualist society that values competency and autonomy Triandis (1989, 1988). The cultural dichotomy between India and the USA (i.e., collectivism vs. individualism) thus presents a rich ground for understanding variations in their moral and ethical values.
Relatedly, gender differences in social behaviors emerge from societal expectations and the process of socialization. For instance, women are less assertive across cultures to avoid negative attribution Amanatullah and Morris (2010); Ferguson and Eyre (2000) and ask for less during negotiation than men Arnold and McAuliffe (2021). We examine how shame, an indicator of negative evaluation, is attributed to women across cultures. We hypothesize that more shame is attributed to women in Indian society due to its patriarchal nature compared to American society.
Movies Total Tokens #shame #pride #control
Bollywood 2738 22.62M 5409 2999 8303
Hollywood 2697 20.78M 1221 2805 4385
5435 43.3M 6583 5804 12688
Table 2: Data Distribution. #shame indicates the number of dialogues with the word "shame" or its derivative form (e.g., ashamed, shameless). Similarly, #pride indicates the number of dialogues with the word pride or its derivative form (e.g., proud). Control is the group of dialogues without words shame, pride, and their derivatives.
2 Background
The vision of safe and accountable AI is centered on LLMs’ moral and value alignment. SOCIAL-CHEM-101 Forbes et al. (2020), SCRUPLES Lourie et al. (2021), SOCIAL BIAS FRAMES Sap et al. (2019), MORAL INTEGRITY CORPUS Ziems et al. (2022), and VALUE PRISM Sorensen et al. (2023) are a few datasets developed to teach socially aligned interactions to LLMs. The dominant approach for norm discovery involves prompting LLMs, sometimes coupled with a verification step such as an entailment test or underlying emotion (negative emotion $\rightarrow$ norm violation) Jiang et al. (2021); Fung et al. (2022); CH-Wang et al. (2023). However, most of the social situations in these datasets and the human annotators employed to label those situations reflect English beliefs and ethics. Additionally, language models prompted to identify "social norms" are known to reflect the values and beliefs of WEIRD nations (Western, Educated, Industrialized, Rich, and Democratic) Havaldar et al. (2023b); Atari et al. (2023). Capturing diversity in morality is thus a research question of great significance Talat et al. (2021).
In this paper, we examine the cross-cultural variations in self-conscious emotions and use these emotions to extract culture-specific normative expectations. We validate our findings by evaluating their alignment with known cultural tendencies of Indian and American societies. Additionally, we examine gender-wise attribution of shame and norm distribution to contextualize the potential impact of inculcating automatically extracted social values into machines.
3 Dataset
Culture-specific people’s conversations are needed to study moral emotions and to mine underlying normative expectations. Social media QnA posts are primarily recollections or ’what-if’ situations that do not reveal fully how these emotions are expressed in natural conversations, for instance between a boss and an employee or among family members at a Christmas table.
Movies are a rich source of real-time social interactions illustrating power dynamics, negotiations, and self-conscious emotions in action; to study how and why these emotions are expressed. We collected English subtitles for 5,435 movies belonging to Bollywood and Hollywood that were released post-1990 by automatically crawling websites that host or link movie subtitles (See Table 2 for movie distribution). The choice of English language for Bollywood movies was made based on the availability of subtitles. The year of release for movies was verified by either parsing subtitle file names having a release year or checking Wikipedia entries. The year mapping was performed to ensure a similar period for collected movies.
3.1 Data Proprocessing
We extracted dialogues having the word shame or pride or their variations (See Appendix A1), along with the previous and the next two lines for situational context (See Table 1). In the case of short dialogues, which could be the case due to monosyllabic responses in spoken conversations, an extra previous and next line was appended to context. One author manually checked the entire dataset of shame and pride-related dialogues to correct spelling mistakes in subtitles, remove unrecognizable characters, and identify phrases having pride and shame-related keywords but not indicating social sanction (or approval).
For instance, the phrases like "what a shame" or "it’s a shame" in Hollywood movies do not reflect social disapproval, and therefore such dialogues were excluded. Below is an example of such dialogue.
‘‘…His music is sad all of the time. It’s really a shame. He’ll never be allowed to go up to heaven.   But you will, won’t you, Johan?   Oh yes. I’ll get there."
Similarly, the phrase "proudly presents" was used in the context of introducing an event or an artist in Hollywood movies. All dialogues with the phrase "proudly presents" were thus removed for further analysis.
We created four sets of dialogues: (a) shame-related dialogues in Bollywood, (b) shame-related dialogues in Hollywood, (c) pride-related dialogues in Bollywood and (d) pride-related dialogues in Hollywood. Additionally, we also formed a control set of dialogues unrelated to shame and pride for both movie industries (See Table 2 for dialogue distribution).
Table 3: Pearson r with $95\%$ confidence interval for top LIWC categories significantly correlated ($p<0.05$) with shame for Bollywood and Hollywood after Benjamini-Hochberg p-value correction. Note the contrast that is, Self-centered shame vs Other-centered shame; Past focus vs Present focus in Hollywood vs Bollywood. See Table A6 and A8 for the complete set of correlations.
4 Approach
Fig. 1 illustrates our study design comprising two key approaches (a) a vocabulary approach to measure cross-cultural variations in linguistic expressions of shame/pride and (b) prompting LLM approach to extract latent normative expectations and examine how they differ between India and the U.S.A.
4.1 Vocabulary Approach: LIWC
Linguistic Inquiry of Word Count (LIWC) Boyd et al. (2022) is a corpus analysis tool widely used in psychology to identify psychological properties of words such as "self-focus" (e.g., I, my, mine) vs. "other-focus" (e.g., you, yours, theirs) which are of great significance for examining social behaviors such as self-regulation and conformity.
To understand cross-cultural linguistic variations in the manifestation of shame and pride, we extracted the normalized distribution of psycho-social categories from the dialogues (shame+control vs pride+control for both movie industries) and examined their correlation with shame and pride. Please note that search keywords (See Table A1) were removed from the LIWC dictionary to prevent overestimation of shame- (e.g., negative emotion) and pride-related categories (e.g., achievement).
4.2 Pretrained LLM: GPT-4
Taking into account pre-trained LLMs’ tendency to favor an Anglocentric understanding of text Havaldar et al. (2023b), we designed the prompts to elicit the "reason behind self-conscious emotions" rather than extracting "social norm" to overcome the cultural bias. The prompts were designed to seek the following:
  • who is $<$being shamed/feeling proud$>$ in the given movie discourse, and what is their gender?
  • What is the reason behind $<$the feeling of shame/pride$>$?
The first question orients the LLM to focus on the subject, i.e., person experiencing the social sanction or approval and then identify their gender. The output for the second question serves as the implicit social norm in the culture. Since GPT-4 exhibits performance at par with human annotators Pan et al. (2023); Törnberg (2023), we designed this task using GPT-4 chat in a two-shot setting. The temperature was set to 00. The prompts with two-shot samples are provided in Tables A4 and A5.
Themes in Social Norms
To understand differences in situations that evoke shame and pride in Indian and American societies, we performed agglomerative clustering after embedding extracted norms using SBERT embeddings Reimers and Gurevych (2019). For this purpose, we merged shame-related norms from Bollywood and Hollywood and pride-related norms from Bollywood and Hollywood and removed duplicates. Social norms extracted from GPT-4 chat are short phrases (See Table 5) and are devoid of culture-specific style markers (i.e., Indian English vs American English) thus, no control for culture is needed.
Table 4: Pearson r with $95\%$ confidence interval for top LIWC categories significantly correlated ($p<0.05$) with pride for Hollywood and Bollywood after Benjamini-Hochberg p-value correction. See Tables A7 and A9 for all significantly correlated categories.
5 Results
5.1 Cross-cultural variations in expressions of shame and pride
Shame
The selected top LIWC categories significantly correlated with shame in Bollywood and Hollywood is provided in Table 3. Although shame is associated with negative emotions, power, and morality in both industries (See Table A6 and A8 for all significant correlations), significant cultural nuances exist in its manifestation. In Hollywood, shame is I-focused with remorse (e.g., sadness, negative emotions, anxiety), whereas in Bollywood, it is You-focused with more social references (e.g., you, your, he, her). Shame is past-focused in Hollywood, whereas it is present-focused in Bollywood movies. Anxiety is unique to Hollywood-shame whereas anger is more strongly associated with shame in Bollywood. Categories such as female, sexual, and swearing are exclusively correlated with shame in Bollywood indicating a conservative society whereas it is illness that is uniquely correlated with shame in Hollywood indicating social sanctions around incapability.
Pride-
related interactions are family focused and self-centered (e.g., I am proud of __) in Hollywood movies whereas it is achievement/power focused and "We"-centered in Bollywood movies (See Table 4). Similarly, pride expressions related to power, morals, and prosocial behavior are more common in Bollywood dialogues, whereas pride expressions related to individuals’ growth (i.e., male/female, work) are more frequent in Hollywood dialogues. Both industries associate pride with ethnicity.
5.2 Cross-cultural variations in Social norms
Manual Evaluation
Two volunteers manually verified the predicted gender for "the person experiencing the self-conscious emotions (shame and pride)" and the reason behind self-conscious emotions in a randomly sampled set of $100$ dialogues from Hollywood and $100$ dialogues from Bollywood. The annotator who labeled the Bollywood set is an Indian, aware of social roles and expectations in Indian society. Likewise, the annotator for the Hollywood set is an American with a nuanced understanding of the social norms of the U.S.A.
Gender was incorrect for eight samples in the Bollywood set and for five samples in the Hollywood set. For Bollywood, GPT-4 chat predicted a gender for 15 (3 predicted as Female, rest as male) samples where gender was not evident, whereas, for 10 (3 Female) such samples in Hollywood. One such example is provided below:
e.g. We’re proud of Bristol’s decision to have her baby… and even prouder to become grandparents? We’re not proud that… our teenage daughter is pregnant, Maria. And I don’t want to send a message that teen pregnancy… is something to be proud of. I want that line out. Of course, Governor.
Here, the gender is predicted as "Male" for Governor. GPT-4 chat returns the most likely gender indicating bias however, it is not always incorrect. It is thus safe to assume that prediction error for gender lies between 8-23% for Bollywood and 5-15% for Hollywood.
For Bollywood, 11 predicted reasons behind self-conscious emotion were incorrect whereas for Hollywood, 2 such samples were incorrect. The reason was not explicit for another nine samples in Bollywood set and one sample in Hollywood. Overall, the prediction error lies between 11-20% for Bollywood whereas it is 2-3% for Hollywood.
Bollywood
eavesdropping on private conversation
expressing love for a man
Incestuous relationship
giving birth to a girl child
Poor academic performance
Hollywood
not living up to expectations
hiding/avoiding confrontation
not returning calls after intimacy
mistreatment of a woman
offering poor quality goods for sale
Table 5: A subset of reasons extracted from movie dialogues expressing shame. A total of 4604 unique reasons (Bollywood-3660, Hollywood-944) were extracted.
Dataset of cross-cultural social norms
For the Bollywood set, GPT-4 chat predicted reasons (also implicit social norms) for 5321 (99.2%) shame-related dialogues out of 5363, and 2237 (74.6%) pride-related dialogues out of 2999. For the Hollywood set, GPT-4 chat predicted a reason for 1156 (94.6%) shame-related dialogues out of 1221 and 1731 (61.7%) pride-related dialogues out of 2805. Upon manual analysis, we found that pride is expressed even without a reason to express affection, specifically toward close family members such as son. GPT-4 chat predicted a total of 10,445 social norms (See Table A2 for norm distribution). A sample of five reasons extracted for shame and pride is provided in Tables 5 and 6.
Bollywood
mastering a trick
fulfilling father’s dreams
provide care for old parents
fiancee’s physical appearance
his wealth
Hollywood
for being a hard worker regardless of the task
being a brilliant student
winning olympic gold
achievements and growth
coming out as queer
Table 6: A subset of reasons extracted from movie dialogues expressing pride. A total of 3163 unique reasons (Bollywood-1589, Hollywood-1574) were extracted.
Figure 2: Relative association ($\Delta$) of Bollywood and Hollywood to themes obtained from agglomerative clustering performed on shame-related norms.
5.2.1 Themes of Social Norms
We found twenty-five clusters for shame-related norms and twelve for pride-related norms using agglomerative clustering ( See Tables A12 and A13 for clustering parameters). The clusters were manually assigned a theme (as depicted on the Y-axis of Fig. 2 and 3) after analyzing the ten closest samples based on cosine distance within the cluster (See Table A10 and A11 for examples). We computed the relative association for each cluster with Bollywood and Hollywood using eq. 1.
$\vec{\Delta}=\forall_{t_{i}\in themes}\frac{\mathcal{D}_{bolly_{t_{i}}}}{{\mathcal{D}_{bolly}}}-\frac{\mathcal{D}_{holly_{t_{i}}}}{\mathcal{D}_{holly}}$ (1)
Here, $\mathcal{D}_{industry_{t_{i}}}$ represents dialogues to $industry\in\{Holly,Bolly\}$ and ${t_{i}}\in themes$ (as depicted on Y-axis).
Shame-related social norms
Poverty and inappropriate behaviors are more strongly associated with shame in Hollywood whereas gender conformity and disrespect are more strongly associated with shame in Bollywood (See Fig. 2). As expected, collectivist factors such as parent-related roles and deviation from family norms are strongly associated with shame in Bollywood whereas individualistic attributes such as incompetence and poor social etiquette evoke shame in Hollywood.
Pride-related social norms
Duty, doing the "right" thing, and, achievement are associated with pride in Hollywood whereas serving the nation, family honor, and son’s success are associated with pride in Bollywood (See Fig. 3).
Figure 3: Relative association ($\Delta$) of Bollywood and Hollywood to themes obtained from agglomerative clustering performed on pride-related norms.
5.2.2 Gender differences in Attribution of Self-conscious Emotions
For the Bollywood set, GPT-4 predicted 1632 targets as female and 4361 as male. For the Hollywood set, GPT-4 predicted 482 targets as female and 1367 as male. Across all combinations (shame vs pride x Bollywood vs. Hollywood in Table A2), there are more male targets than females. This aligns with Geena Davis Inclusion Quotient (2016) findings that male characters get more screen time and speak more than female characters.
Using target gender as predicted by GPT-4 (See Table A2 for distribution), we computed the gender-wise attribution to emotions "shame" and "pride" using eq. 2. A positive score indicates a higher association of gender groups with pride, whereas a negative score reflects a higher association with shame. A null score indicates no preference.
$\vec{\Delta}_{g}=\forall_{g\in\{male,female\}}\frac{\mathcal{D}_{pride_{g}}-\mathcal{D}_{shame_{g}}}{\mathcal{D}_{g}}$ (2)
Here, $\mathcal{D}_{pride_{g}}$ represents movie dialogues having pride-related word where the target of pride is $g\in\{male,female\}$, $\mathcal{D}_{shame_{g}}$ represents movie dialogues having shame-related word where the target of shame is $g\in\{male,female\}$ and $\mathcal{D}_{g}$ represents total movie dialogues grouped based on gender. As depicted in Fig. 4, Hollywood movies are pride-oriented, whereas Bollywood movies are shame-oriented. Females are attributed more shame, and the difference (male-female) in the expression of pride and shame is similar (0.16 for Hollywood and 0.18 for Bollywood) irrespective of the movie industry.
Normative expectations evoking shame are surprisingly similar for females across movie industries (See Figs. 5 and 6). Sexuality (e.g., promiscuity, immodesty, gender roles) is the dominant theme for women in both movie industries whereas males are shamed for incompetency in both industries. For pride, women express more pride for family roles. However, we do see a variation in themes for men-pride i.e., justice, bravery and achievement are major themes for pride in Bollywood whereas it is duty, and self-identity in Hollywood. We also note interesting reverse patterns e.g., self-identity is a source of pride for females in Bollywood whereas it is for males in Hollywood.
Figure 4: Relative association ($\Delta$) of emotions pride and shame gender-wise in Hollywood and Bollywood dialogues. A higher positive score indicates a stronger association of gender with pride.
6 Discussion
Cross-cultural variations in expressions of moral emotions
The entailment of shame tells us about undesired behavior, whereas pride reveals desired behaviors in a society. While self-conscious emotions are culture-agnostic markers of unspoken social rules, the expressions of these emotions vary across cultures. We uncovered these variations in Sec 5.1. Shame is a highly undesirable self-focused emotion in the U.S. emphasizing incompetency and failures and is rarely used Cohen (2003); Boiger et al. (2013). Notably, we also observe a strong correlation of first-person singular pronouns with shame and its infrequent use in Hollywood movies compared to Bollywood movies (See Table 2). In contrast, shame is interdependent in collectivist communities Wong and Tsai (2007). Our empirical analyses reveal You-focused (e.g. You, social references) nature of shame in Bollywood movies (See Table A6). Moreover, the contrasting tenses coupled with different emotions, i.e., past + sadness in Hollywood vs. present + anger in Bollywood, reflect their varying goals, i.e., remorse for past failures/losses vs enforcing conformity Wong and Tsai (2007). The correlation of shame with categories such as sexual, female, conflict, swear is exclusive to Bollywood, likely due to emphasis on conformity in Indian society. Pride-related discourse in Hollywood is duty and achievement-focused, in line with prior findings underlining the increasing significance of "success" with growing individualism Cohen (2003), whereas in Bollywood, pride is centered around collective achievement (we, achievement, prosocial).
Moral Emotion-based Prompting
Compared to GoEmotions dataset Demszky et al. (2020) having 817 samples for embarrassment and 452 samples for pride, we provide a much larger cross-cultural dataset of shame (6583 samples) and pride (5804 samples) related discourse, collected from movie subtitles across two cultures. Additionally, we release 10K culture-specific social norms. These can be used for downstream applications like evaluation of culturally aware language inference and achieving pluralistic values in LLMs.
Self-conscious emotion-based prompting helped in mitigating cultural bias in LLMs-generated text and potential stereotyping during norm discovery; as evident from the specificity of the norms extracted using our approach ( See Tables 5 and 6). More importantly, these norms reflect the cultural subtleties (e.g., desire for a son, parents’ duty to marry off their daughter vs. honesty in business, returning calls after date night as in Table 5) revealing the differences in socially expected behaviors in Indian society and the U.S. society. The excerpts explicitly expressing shame and pride from movie subtitles may have led to the over-representation of certain social situations (e.g., son’s achievement and daughter’s wedding in India vs duty and competence in America). A natural next step would be to capture situations with implicit shame and pride. The pipeline in Fig. 1 could be adapted to first detect self-conscious emotion in a given discourse and then identify the reason for advanced computational cultural understanding of the text. For this purpose, our dataset of shame/pride-related situations will be useful.
Cross-cultural variations in Social Norms
To tease apart underlying values evoking moral emotions across cultures, we first considered mapping social norms to Schwartz’s Theory of Values Schwartz (2012). However, a social norm can have contrasting values depending on the culture. Consider "refusing to marry", an instance of non-conformity in Indian society, whereas it is an instance of self-direction in a Western context. Of course, society is evolving and we acknowledge that urbanized regions of India may also consider this an instance of self-direction. However, the diachronic shift in the underlying values of social norms is out of the scope of this paper. To effectively understand the cross-cultural variations in social norms, we performed hierarchical clustering and empirically picked the distance after manually analyzing the quality and granularity of clusters. The situations (i.e., poverty, inappropriate behavior, incompetence) connected to shame in the U.S. society vs situations(i.e., gender roles, disrespect, family norms) in Indian society as evident in Fig 2 align with known cultural tendencies of both nations. Likewise, pride is strongly associated with duty and achievement in Hollywood whereas with collective achievement (i.e., nation honor, family honor) in Bollywood. While the clustering was performed on the reasons extracted from movie conversations, the cluster themes still reflect the patterns seen during LIWC analysis, supporting the use of moral emotions for overcoming cultural bias in LLMs.
Gender Differences
Shame, a self-conscious emotion expressing devaluation, is targeted towards women more than men, whereas pride, a positive valence emotion expressing value, is used for more men than women in movies. Surprisingly, the gender gap in pride- and shame-related dialogues is similar across both societies. While our initial hypothesis that "more shame is attributed to women in Indian society compared to American society" (0.18 in Bollywood vs 0.16 in Hollywood) is true, the difference is negligible. It is also surprising to see the similarity in normative expectations associated with genders male and female in Indian and American societies. Our analysis reveals that women are more frequent targets of social sanctions than men. It is thus important to characterize the source and acquired social norms before their use for achieving pluralistic values in LLMs otherwise we risk social bias.
7 Conclusion
We introduce an emotion-based approach to extract and analyze social norms across cultures. We create the largest known multi-cultural dataset of self-conscious emotions and relevant underlying norms. Our post hoc analysis of social norms demonstrates (a) cross-cultural linguistic differences in self-conscious emotions (shame and pride), (b) cultural dichotomy in normative expectations in India and the U.S.A., and (c) more "social sanctions" and fewer "social benefits" to the female gender in movies.
Future work can utilize our dataset to better align LLMs to various cultures. Additionally, our analysis of self-conscious emotions is the first of its kind; we hope future NLP researchers will build upon this work to investigate social norms in LLMs from a multicultural, emotion-based perspective.
Social Impact and Ethics Statement
Social norms discovery is a crucial component in programs [^2] ALIGN-https://www.alignplatform.org/learning-collaborative designed for social and behavioral change to promote equity, social justice, and well-being Mauduy et al. (2022); Bonan et al. (2020). Further work explores style as a product of norm differences Havaldar et al. (2023a). Social psychology investigates social norms (descriptive vs injunctive) to design experiments for understanding behaviors such as self-regulation, persuasion Cialdini et al. (1990) and decision-making Gavrilets (2020); Bhanot (2021) to promote collective-level change in societies.
Relatedly, Kimbrough and Vostroknutov (2023) showed people’s tendency to choose self-serving social norms using a dictator-recipient setup, emphasizing the need for dedicated research efforts to understand morality and belief distortion in different contexts. The norms and cultural preferences learned from movies that often showcase stereotypical behaviors of society may induce pluralistic ignorance and, more importantly, lead to discrimination and biases in LLMs when used for training. We hope that this paper will encourage scrutiny of source corpora and derived norms before their use for fine-tuning LLMs.
Limitations
Social norms mutate as society evolves. We acknowledge that our dataset of movies (released post-1990) may reflect social norms that are less characteristic of contemporary society. Moreover, countries like India and America contain a mix of cultures. The captured norms may not reflect the cultural variations, for example, between regions (e.g., East Coast vs West Coast in the U.S.A or North India vs South India). Movies also exaggeratedly depict the world around us (e.g., weddings, criminal activities, sexual abuse, etc.), and we caution against stereotyping cultures based on movie-based norms.
The dominant language in Bollywood movies is Hindi and our analysis is based on their English translations which may not always be accurate, especially when the discourse is about concepts native to a culture. Relatedly, the LIWC may not have high coverage for such concepts. We did not compare the movie genre and acknowledge that situational/unrealistic norms (e.g., science fiction, comedy, etc.) could exist. Regardless, social norms associated with shame and pride are still relatable and reflect the target audience’s beliefs. It is also worth noting that gender norms in movies may not fully reflect the norms of today’s society. This study is conducted over a thirty-year long period and thus, the aggregated norms may not reflect the current trends.
The prompts ("You study Indian society" vs "You study Western society" described in Tables A4 and A5 ) used to elicit norms may have induced unwanted bias Cheng et al. (2023); Lucy and Bamman (2021) and it is worth investigating the variations if any, in extracted norms due to different prompt designs.
Appendix A Appendix
A.1 Search Keywords
Search keywords defined in Table A1 were used to extract shame- and pride-related dialogues from Bollywood and Hollywood movie subtitles.
Lemma Word forms
shame shame, shamed, shameful, ashamed
pride proud, prouder, proudly, pride
Table A1: List of search keywords for extracting dialogues.
A.2 Norms Distributions
Over 10k reasons (or norms) were extracted using GPT-4. The distribution of norms and gender of targets is provided in Table A2.
Bollywood Hollywood
Gender shame pride shame pride
male 3102 1259 591 776
female 1306 326 246 236
not known 913 652 319 719
total 5321 2237 1156 1731
Table A2: Norm and Gender distribution (with duplicates) for dialogues for which GPT-4 predicted male or female. The duplicate norms are not removed as their frequency reflect their prevalence and is useful for estimating gender association.
A.3 Annotation
The annotation guidelines to verify the gender predicted by GPT-4 and the correctness of the reason are provided in Table A3. The annotators for Bollywood set and Hollywood set were Indian and American respectively. Both annotators were female, proficient in English language, and well-versed with social norms. During annotation, if the gender or the reason is unclear, the annotators were asked to label "not explicit". We only considered the cases where the gender was predicted to be either male or female. The task is objective and inter-annotator agreement was not computed. The annotators volunteered for the task and were not provided monetary compensation.
Guidelines for Manual Evaluation
1. Read the conversation and identify the person feeling ashamed (or being shamed) or proud.
2. Identify the gender. Check gender markers such as Mr/Mrs., s/he, him/her, etc. If the name is provided in the conversation, check if the name is likely to be a male name or female. If not clear, mark "not explicit".
3. Read the reason behind shame/pride. Compare with conversation and determine if the provided reason is the cause for shame/pride.
Table A3: Guidelines for Annotation
A.4 Prompts
The prompts designed to extract the subject experiencing self-cosncious emotions, their gender and the reason behind the emotions are provided in Tables A4 and A5.
Prompts for norm discovery in Bollywood movie dialogues
"You are an experienced social scientist and you study Indian society. Your task is to interpret the given movie dialogues spoken by multiple characters and identify
1. Who is feeling proud? Provide the gender. If agent or gender is not explicit, use the pronouns, name, and salutations to guess agent and gender.
2. Identify the reason behind feeling proud. If not explicit, return NA.
The answer should be short and in below CSV format. $<$who_proud, gender, reason$>$
Input: Your mother had already given up on me because I was useless hopeless But you were the apple of her eye My child will make me proud look after me He’ll support in my old-age But this kid left her all alone.
Output: mother, male, provide care for old parents
Input: I’ve heard there’s a promising young student in your school What’s his name? He’s made us proud in long jump, we are here to felicitate him Call him Show yourself, Raju Tempre
Output: authority, NA, Sports achievement"
"You are an experienced social scientist and you study Indian society. Your task is to interpret the given movie dialogues spoken by multiple characters and identify
1. Who is being shamed? Predict the gender. If gender is not explicit, use the pronouns, name, and salutations to guess gender.
2. Identify the primary reason for shaming
The answer should be short and in CSV format. $<$ shamed, gender, reason$>$
Input: And should we bow before others begging....them to marry our daughters? This shall not happen. Neither will the girls be alive here nor shall....we be ashamed of ourselves. You cannot kill the life which God has given. I won’t let you commit the sin.
Output: girl’s parent, NA, not able to marry off their daughters
Input: Black marketers are now in the open. And the thieves too Politics is in a great mess Shame on this system. There’s no democracy Get rid of these politicians The gong has struck..’’Our hearts are swaying to it’s beats’’
Output: System, NA, poor law and regulations"
Table A4: Bollywood: Prompts for norm discovery using GPT-4 Chat.
Prompts for norm discovery in Hollywood movie dialogues
"You are an experienced social scientist and you study Western society. Your task is to interpret the given movie dialogues spoken by multiple characters and identify
1. Who is feeling proud? Provide the gender. If agent or gender is not explicit, use the pronouns, name, and salutations to guess agent and gender.
2. Identify the reason behind feeling proud? If not explicit, return NA.
The answer should be short and in below CSV format. $<$who_proud, gender, reason$>$
Input: I want to go to Worlds and win gold. I want to go to the 88 Olympics in Seoul and win gold. Good! I’m proud of you. Are you getting the support that you need? What do you mean sir?
Output: Sir, male, winning olympic gold
Input: Yes. Yes, I did. I promise, this time I really got the promotion. - I’m proud of you, son. - Thank you, sir. Excuse me. Hi, sweetheart.
Output: father, male, for getting the promotion"
"You are an experienced social scientist and you study western societies. Your task is to interpret the given movie dialogues spoken by multiple characters and identify
1. Who is being shamed? Predict the gender. If gender is not explicit, use the pronouns, name, and salutations to guess gender.
2. Identify the primary reason for shaming.
The answer should be short and in CSV format. $<$ shamed, gender, reason, prevalence $>$
Input: You still owe me 100. Remember? You stiffed Donny for 100 bucks? Cheapskate. Shame on you. Pay this man his C-note. Now I know why they call you the Snake.
Output: NA, male, not returning borrowed money
Input: You prey on your own people. You steal from your own people. Have you no shame!? - Huh? - Well, we’re still here. Man: Mr. Markopolos, it’s all yours.
Output: Snake, male, stealing and preying on people"
Table A5: Hollywood: Prompts for norm discovery using GPT-4 Chat.
A.5 LIWC Correlation Results
Tables A6, A7, A8 and A9 contain the significantly correlated ($p<0.05$) LIWC categories, the most frequent five words for each category, pearson r and 95% confidence interval.
LIWC Categories Top-5 words correlation 95% CI
EMO_NEG (bad, mad, scared, worry, fear) 0.330 [0.315, 0.344]
TONE_NEG (lost, kill, wrong, bad, hit) 0.249 [0.233, 0.265]
POWER (sir, respect, own, kill, poor) 0.198 [0.182, 0.214]
EMOTION (love, good, bad, happy, crazy) 0.168 [0.152, 0.185]
YOU (you, your, you’re, yourself, you’ve) 0.161 [0.145, 0.178]
SOCREFS (you, your, he, her, him) 0.148 [0.132, 0.165]
FEELING (feel, touch, feeling, felt, hard) 0.133 [0.117, 0.150]
DRIVES (we, our, us, sir, married) 0.115 [0.098, 0.132]
SOCIAL (you, your, he, her, him) 0.110 [0.094, 0.127]
MORAL (wrong, innocent, duty, decent, excuse) 0.102 [0.085, 0.118]
AFFECT (love, good, keep, respect, well) 0.078 [0.062, 0.095]
NEGATE (not, don’t, no, aren’t, won’t) 0.065 [0.049, 0.082]
FEMALE (her, she, girl, she’s, mom) 0.060 [0.043, 0.077]
PPRON (you, i, me, your, my) 0.058 [0.041, 0.075]
FAMILY (son, married, uncle, dad, mom) 0.055 [0.039, 0.072]
PREP (to, of, in, for, on) 0.053 [0.037, 0.070]
SEXUAL (chaste, lust, sex, sexy, pimp) 0.051 [0.034, 0.067]
PRONOUN (you, i, me, your, my) 0.040 [0.023, 0.057]
AUXVERB (is, are, have, be, don’t) 0.029 [0.013, 0.046]
CONFLICT (kill, killed, accusing, killing, cruel) 0.024 [0.007, 0.041]
SWEAR (hell, bloody, idiot, damn, ass) 0.024 [0.007, 0.041]
EMO_ANGER (mad, angry, hate, cruel, argue) 0.022 [0.005, 0.039]
FOCUSPRESENT (is, are, don’t, i’m, aren’t) 0.021 [0.005, 0.038]
Table A6: Psychosocial categories significantly correlated (p<0.05) with shame in Bollywood dialogues. p-values were corrected using Benjamini-Hochberg correction. The categories are arranged in decreasing order of correlation.
LIWC Categories Top-5 words correlation 95% CI
DRIVES (we, our, us, sir, work) 0.122 [0.104,0.140]
DET (the, a, my, your, that) 0.101 [0.082,0.119]
ACHIEVE (work, better, win, best, try) 0.098 [0.080,0.116]
POWER (sir, own, respect, kill, power) 0.091 [0.072,0.109]
SOCREFS (you, your, he, we, our) 0.085 [0.067,0.103]
PREP (of, to, in, for, with) 0.084 [0.066,0.103]
MORAL (wrong, duty, brave, arrogant, useless) 0.075 [0.057,0.093]
CONJ (and, but, so, if, as) 0.075 [0.056,0.093]
REWARD (win, won, glory, success, successful) 0.071 [0.052,0.089]
TONE _POS (love, good, thank, well, great) 0.064 [0.046,0.082]
POLITIC (nation, army, sultan, president, dynasty) 0.064 [0.045,0.082]
WE (we, our, us, we’ll, let’s) 0.060 [0.042,0.079]
SOCIAL (you, your, he, we, our) 0.060 [0.041,0.078]
FAMILY (son, papa, married, dad, uncle) 0.059 [0.041,0.078]
AFFILIATION (we, our, us, dear, we’ll) 0.059 [0.041,0.077]
FEELING (feel, feeling, hard, felt, sense) 0.055 [0.036,0.073]
ETHNICITY (indian, indians, british, hindi, caste) 0.054 [0.036,0.072]
MALE (he, his, him, son, sir) 0.054 [0.036,0.072]
CULTURE (indian, nation, army, car, indians) 0.044 [0.026,0.062]
AFFECT (love, good, thank, well, great) 0.040 [0.021,0.058]
ARTICLE (the, a, an, tha) 0.039 [0.021,0.058]
PPRON (you, i, my, your, me) 0.036 [0.017,0.054]
PROSOCIAL (thank, please, sorry, respect, gift) 0.032 [0.014,0.051]
FUNCTION (you, the, i, of, to) 0.032 [0.013,0.050]
YOU (you, your, you’re, you’ve, you’ll) 0.030 [0.011,0.048]
THEY (they, their, them, they’re, they’ll) 0.029 [0.011,0.048]
CERTITUDE (really, real, surely, proved, actually) 0.021 [0.002,0.039]
EMO _POS (love, good, happy, happiness, smile) 0.020 [0.001,0.038]
Table A7: Psychosocial categories significantly correlated (p<0.05) with pride in Bollywood dialogues. p-values were corrected using Benjamini-Hochberg correction. The categories are arranged in decreasing order of correlation.
LIWC Categories Top-5 words correlation 95% CI
EMO_NEG (sick, pain, fear, bad, afraid) 0.425 [0.403, 0.446]
TONE_NEG (lost, wrong, sick, pain, poor) 0.355 [0.331, 0.377]
EMOTION (good, love, sick, pain, bad) 0.290 [0.266, 0.314]
POWER (own, sir, poor, killed, war) 0.263 [0.238, 0.287]
AFFECT (well, good, love, help, damn) 0.168 [0.142, 0.193]
DRIVES (we, us, our, work, we’re) 0.152 [0.127, 0.178]
FUNCTION (you, i, the, to, of) 0.131 [0.105, 0.157]
PPRON (you, i, me, i’m, my) 0.111 [0.085, 0.137]
MORAL (wrong, excuse, decent, honest, duty) 0.110 [0.084, 0.136]
I (i, me, i’m, my, i’ll) 0.102 [0.076, 0.128]
EMO_SAD (crying, cry, sob, lonely, sad) 0.096 [0.070, 0.122]
NEGATE (no, not, don’t, nothing, never) 0.090 [0.063, 0.116]
PREP (to, of, in, for, on) 0.089 [0.062, 0.115]
PRONOUN (you, i, that, it, me) 0.088 [0.062, 0.114]
YOU (you, your, you’re, yourself, you’ve) 0.080 [0.054, 0.106]
AUXVERB (be, i’m, is, was, have) 0.077 [0.051, 0.103]
FOCUSPAST (was, did, were, been, didn’t) 0.069 [0.043, 0.096]
SOCIAL (you, your, we, he, you’re) 0.060 [0.033, 0.086]
CONJ (and, so, but, if, when) 0.056 [0.030, 0.082]
LINGUISTIC (you, i, the, to, of) 0.056 [0.030, 0.082]
ALLNONE (no, all, nothing, never, yes) 0.055 [0.029, 0.081]
FAMILY (son, dad, baby, mom, mama) 0.054 [0.028, 0.081]
SOCREFS (you, your, we, he, you’re) 0.050 [0.024, 0.077]
EMO_ANX (fear, afraid, worry, terrified, scared) 0.042 [0.015, 0.068]
ILLNESS (sick, pain, pains, flu, sickly) 0.038 [0.012, 0.064]
FEELING (feel, felt, pain, feeling, hard) 0.035 [0.008, 0.061]
EMO_ANGER (hate, hated, mad, angry, hates) 0.033 [0.007, 0.059]
DIFFER (not, but, if, didn’t, or) 0.033 [0.007, 0.059]
DISCREP (should, can, would, can’t, want) 0.032 [0.006, 0.059]
COGNITION (no, not, all, know, but) 0.032 [0.005, 0.058]
Table A8: Psychosocial categories significantly correlated (p<0.05) with shame in Hollywood dialogues. p-values were corrected using Benjamini-Hochberg correction. The categories are arranged in decreasing order of correlation.
LIWC Categories Words correlation 95% CI
PPRON (you, i, i’m, me, my) 0.135 [0.113,0.158]
SOCREFS (you, your, we, he, you’re) 0.131 [0.108,0.154]
FAMILY (son, dad, baby, mom, mama) 0.121 [0.099,0.144]
CONJ (and, so, but, as, if) 0.114 [0.091,0.137]
SOCIAL (you, your, we, he, you’re) 0.102 [0.079,0.125]
FUNCTION (you, i, the, of, to) 0.100 [0.077,0.123]
I (i, i’m, me, my, i’ll) 0.095 [0.072,0.118]
YOU (you, your, you’re, you’ve, yourself) 0.091 [0.068,0.114]
MALE (he, his, him, man, son) 0.084 [0.061,0.107]
DRIVES (we, our, us, we’re, dad) 0.076 [0.053,0.099]
TONE _POS (good, well, thank, great, love) 0.072 [0.049,0.095]
PRONOUN (you, i, i’m, that, it) 0.072 [0.048,0.095]
AUXVERB (i’m, be, is, was, have) 0.071 [0.048,0.094]
EMO _POS (good, love, happy, hope, wonderful) 0.070 [0.047,0.093]
PREP (of, to, in, for, on) 0.070 [0.047,0.093]
AFFILIATION (we, our, us, we’re, dad) 0.054 [0.031,0.078]
EMOTION (good, love, happy, hope, bad) 0.052 [0.029,0.075]
ETHNICITY (american, irish, chinese, german, christian) 0.052 [0.029,0.075]
REWARD (win, won, winner, successful, earned) 0.048 [0.025,0.071]
ACHIEVE (work, better, best, trying, try) 0.045 [0.022,0.068]
POWER (sir, own, war, strong, mighty) 0.044 [0.021,0.067]
AFFECT (good, well, thank, great, love) 0.041 [0.018,0.064]
MORAL (wrong, excuse, hero, brave, dignity) 0.040 [0.017,0.064]
FEMALE (her, she, she’s, girl, ladies) 0.034 [0.011,0.057]
CULTURE (american, car, president, nation, mayor) 0.033 [0.009,0.056]
FOCUSPAST (was, did, been, were, had) 0.032 [0.009,0.055]
SHEHE (he, his, him, her, she) 0.031 [0.008,0.054]
WORK (work, job, school, deal, company) 0.028 [0.005,0.051]
Table A9: Psychosocial categories significantly correlated (p<0.05) with pride in Hollywood dialogues with 95% confidence intervals. p-values were corrected using Benjamini-Hochberg correction. The categories are arranged in decreasing order of correlation.
A.6 Clustering Results
Tables A12 and A13 contain the manually annotated Cluster Theme, the total number of samples in each cluster, and Bollywood vs Hollywood distribution. The distance was set to $5$ and the duplicates were removed. The theme and top three examples demonstrating its meaning are provided in Tables A10 and A11.
Table A10: Themes and top-3 examples for clusters obtained after agglomerative clustering of shame-related dialogues from Bollywood and Hollywood.
Table A11: Themes and top-3 examples for clusters obtained after agglomerative clustering of pride-related dialogues from Bollywood and Hollywood.
Table A12: Distribution of reasons (shame) across manually labeled clusters. A total of twenty-six clusters were generated with distance=5. Duplicates were removed for clustering. One cluster had generic reasons ( indicating lack of shame) which was removed for further analysis. Finally, 25 clusters were considered.
Table A13: Distribution of reasons (pride) across manually labeled clusters. A total of fifteen clusters were generated with distance=5. Duplicates were removed for clustering. One cluster had generic statements (without explicit reason) and was removed for further analysis. ’Achievement’ and ’Winning’ were merged to form ’Achievement’ and, ’Bravery’ and ’Nation’ were merged to form ’Bravery & Nation’. Finally, 12 clusters were considered for analysis.
(a) Bollywood
(b) Hollywood
Figure 5: Gender-wise differences in themes associated with normative expectations from shame-related discourse in Bollywood and Hollywood. $\Delta(Female,Male)$ is the difference between normalized dialogues attributed to females under a theme and to males under the same theme (as in eq. 1 but for dimension gender). Promiscuity is most associated with the female gender in both industries.
(a) Bollywood
(b) Hollywood
Figure 6: Gender-wise differences in themes associated with normative expectations from pride-related discourse in Bollywood and Hollywood. $\Delta(Female,Male)$ is the difference between normalized dialogues attributed to females under a theme and to males under the same theme (as in eq. 1 but for dimension gender). Promiscuity is most associated with the female gender in both industries.