### v3

# Social Norms in Cinema: A Cross-Cultural Analysis of Shame, Pride and Prejudice

Abstract
Social emotions such as shame and pride reflect social sanctions or approvals in society. In this paper, we examine how expressions of shame and pride vary across cultures and harness them to extract unspoken normative expectations across cultures. We introduce the first cross-cultural shame/pride emotions movie dialogue dataset, obtained from  5.4K Bollywood and Hollywood movies, along with over $10K$ implicit social norms. Our study reveals variations in expressions of social emotions and social norms that align with known cultural tendencies observed in the United States and India -- e.g., Hollywood movies express shame predominantly toward self whereas Bollywood movies express shame predominantly toward others. Similarly, Bollywood shames non-conformity in gender roles, and takes pride in collective identity, while Hollywood shames lack of accountability, and takes pride in ethical behavior. More importantly, women face more prejudice across cultures and are sanctioned for similar social norms.
Social Norms in Cinema: A Cross-Cultural Analysis of Shame, Pride and Prejudice
1 Introduction
Normative expectations [^1] Bicchieri (2005) uses the term normative expectations, Lahti and Weinstein (2005) use moral norms, and many social psychologists Cialdini et al. (1990) use injunctive norms. We use the terms social norms and normative expectations interchangeably., a category of social norms that refer to others’ beliefs about how we ought to behave Bicchieri (2005). It reveals attitudes and values that influence social actions across cultures. For instance, hesitancy toward seeking mental help could exist because others expect an individual not to seek help or because the individual believes that others do not seek help. The former reveals the attitude of society, whereas the latter indicates an individual’ lack of awareness. Capturing diversity in cultural beliefs is paramount to building socially aligned AI models for diverse users Talat et al. (2021); Atari et al. (2023), for inducing collective behavior change (e.g., overcoming stigma against mental health disorders) and several other social applications.
Look, Alok is in love with me and I love him too.
Shameless one! Have you lost all your shame?
What is there to be ashamed in this? Don’t sisters love their brothers in this vicinity?
$\Rightarrow$ Incestuous relationship evoke shame.
Sister-in-law! Congrats, sister-in-law! Big brother has started working!
Really?
Now you will have a place of pride in this family!
Yes, please!
$\Rightarrow$ Employed husband evokes pride
Table 1: Excerpts of dialogues expressing explicit social emotions shame and pride indicating actions and social approval.
Prior works crowd-sourced social norms from Q&A style social media posts (such as Reddit) Jiang et al. (2021); Nahian et al. (2020) or prompted large language models (LLMs) Fung et al. (2022) without distinguishing between normative and empirical expectations. However, social media platforms [^2] Twitter Statistic: https://blog.hootsuite.com/twitter-statistics/, [^3] Reddit Statistic: https://www.enterpriseappstoday.com/stats/reddit-statistics.html are skewed toward Western demographics (i.e., young white male) and the reliance on a small cohort of annotators typically located in Western countries further enforces Anglocentric beliefs when collecting socially acceptable behaviors Ziems et al. (2022). LLMs with sophisticated capabilities such as GPT-4 are increasingly used to extract social norms; however, Havaldar et al. (2023a) illustrated their lack of cross-cultural understanding.
Figure 1: An overview of our approach comprising two key steps (a) Vocabulary approach and (b) Prompting a pre-trained LLM.
The overarching question we ask in this paper is, how and what social behaviors are sanctioned or encouraged across cultures. Shame and pride are social emotions experienced across cultures on violating or adhering to normative expectations in society Tracy and Robins (2004). Notably, the function of shame (i.e., social regulation) is consistent across cultures Goetz and Keltner (2007); Fessler (2007); Schaumberg and Skowronek (2022). Examining situations evoking shame and pride can thus reveal unspoken social norms specific to a culture (See Table 1).
In this paper, we (a) examine cross-cultural variations in linguistic expressions of shame and pride, and (b) propose a cross-cultural social emotion-based approach for norm discovery. The appraisals and expressions of shame and pride vary across cultures and is a research question of great significance to understand society and self-regulation Wong and Tsai (2007); Lewis et al. (2010); Wegge et al. (2022). We chose two top-grossing movie industries, i.e., Hollywood and Bollywood, to analyze how social emotions and norms differ across cultures. Hollywood primarily depicts social situations from the USA, an individualist society that values competency and autonomy Triandis (1989, 1988) whereas Bollywood depicts India, a collectivist society where one’s sense of self is interwoven with community beliefs. The cultural dichotomy between India and the USA (i.e., collectivism vs. individualism) thus presents a rich ground for understanding variations in their beliefs and values.
We release a large cross-cultural dataset of social emotions-related dialogues with $10K$ social norms. Using a blend of psychology-informed language analysis and LLM-powered approach to mine implicit information, we demonstrate cross-cultural variations in how and why these emotions are expressed. In parallel, we examine how social sanctions are attributed to men and women across cultures influencing their social behaviors; women are less assertive across cultures to avoid negative attribution Amanatullah and Morris (2010); Ferguson and Eyre (2000) and ask for less during negotiation than men Arnold and McAuliffe (2021). We hypothesize that women are more subjected to social sanctions than men across cultures.
2 Movies Subtitles Corpora
Movies provide life-like depictions of social situations to study how and why these emotions are expressed. They are influenced by and/or, known to influence real lives Adkins and Castle (2014); Kubrak (2020). In contrast, social media Q&A posts are (a) skewed to social situations specific to tech-savvy young demographic groups, and (b) are second-hand reports that do not reveal fully the social dynamics of how these emotions are expressed in natural conversations, for instance, between a boss and an employee or among family members at a Christmas table.
For our study, we collected English subtitles for 5,435 Hollywood and Bollywood movies that were released post-1990 by auto-crawling websites that host or link movie subtitles (See Table A2 for data distribution). The year of release for movies was verified by either parsing subtitle file names having a release year or checking Wikipedia entries. The year mapping was performed to ensure a similar period for collected movies.
2.1 Extracting Shame and Pride related Expressions
The expressions of shame and pride could be implicit (as indicated by lowered shoulders, avoiding eye contact, etc.), and explicit (See Table 1). Detecting implicit social emotions in textual discourse is an innately complex task requiring socio-cultural knowledge that may not be evident to outsiders or third-party annotators. Below is one such example:
‘‘Actor 1: Liza, Alice! Come away from that window this minute. I think it’s disgusting, standing there where they can see you.
Actors 2,3 (Liza, Alice): I notice that you peek out often enough when you think no one’s looking.
Actor 1: I never!
Actors 2,3 (Liza, Alice): I saw you! So did I.
Here, it is difficult to determine if shame is evoked and/or experienced by actors in this discourse. A third-party annotation will reveal their beliefs and attitudes, which could introduce bias in our analysis. Adopting a high precision-low recall strategy, we focus on explicit mentions of shame and pride to learn about occurrences of shame and pride as determined by the characters. We extracted dialogues having the word shame or pride or their variations (See Appendix A1), along with the previous and the next two lines for situational context (See Table 1). In the case of short dialogues, which could be the case due to monosyllabic responses in spoken conversations, an extra previous and next line was appended to context.
Manual Screening One author manually checked the entire dataset of shame and pride-related dialogues to filter out dialogues unrelated to social sanctions/approvals e.g., phrases such as "what a shame", "it’s a shame", "proudly presents". Below is an example of such dialogue.
‘‘…His music is sad all of the time. It’s really a shame. He’ll never be allowed to go up to heaven.   But you will, won’t you, Johan?   Oh yes. I’ll get there."
Similarly, the phrase "proudly presents" is used while introducing an event or an artist in Hollywood movies.
Final Dataset: We created four sets of dialogues: (a) shame-related dialogues in Bollywood, (b) shame-related dialogues in Hollywood, (c) pride-related dialogues in Bollywood and (d) pride-related dialogues in Hollywood. Cross-cultural differences also exist in speaking styles across cultures and we thus formed a control set of dialogues unrelated to shame and pride for both movie industries to remove the variations in language markers owing to culture-specific speaking styles (See Table A2 for dialogue distribution).
Hollywood: LIWC Category (Top-5 words) $r$ Bollywood: LIWC Category (Top-5 words) $r$
Moral (wrong, excuse, decent, honest, duty) 0.110 $2^{nd}$ person pronouns (you, your, you’re, yourself, you’ve) 0.161
$1^{st}$ person sing. pronouns (i, me, i’m, my, i’ll) 0.102 Social References (you, your, he, her, him ) 0.148
Sadness (crying, cry, sob, lonely, sad) 0.096 Moral (wrong, innocent, duty, decent, excuse) 0.102
$2^{nd}$ person pronouns (you, your, you’re, yourself, you’ve) 0.080 Female ( her, she, girl, she’s, mom) 0.060
Past focus (was, did, were, been, didn’t ) 0.069 Family (son, married, uncle, dad, mom) 0.055
All or none (no, all, nothing, never, yes) 0.055 Sexual (chaste, lust, sex, sexy, pimp) 0.051
Family (son, dad, baby, mom, mama) 0.054 Conflict (kill, killed, accusing, killing, cruel) 0.024
Social References (you, your, we, he, you’re) 0.050 Swear words (hell, bloody, idiot, damn, ass) 0.024
Anxiety (fear, afraid, worry, terrified, scared) 0.042 Anger (mad, angry, hate, cruel, argue) 0.022
Illness (sick, pain, pains, flu, sickly) 0.038 Present focus (is, are, don’t, i’m, aren’t) 0.021
Table 2: Pearson r for LIWC categories significantly correlated ($p<0.05$) with shame for Hollywood and Bollywood after Benjamini-Hochberg p-value correction (See Appendix for confidence intervals and p values). Note the contrast that is, Self-centered shame vs Other-centered shame; Past focus vs Present focus in Hollywood vs Bollywood. See Table A7 and A9 for the complete set of correlations.
3 Approach
Fig. 1 illustrates our study design comprising a blend of two key approaches (a) a psychosocial vocabulary approach to measure how expressions of shame and pride differ and (b) prompting Large Language Model approach to extract reasons behind them and how they differ between India and the U.S.A.
3.1 Vocabulary Approach: LIWC
Linguistic Inquiry of Word Count (LIWC) Boyd et al. (2022) is a corpus analysis tool widely used in psychology to identify psychological categories of words such as "self-focus" (e.g., I, my, mine) and "other-focus" (e.g., you, yours, theirs) which are of great significance for examining social behaviors such as self-regulation and conformity.
To understand cross-cultural linguistic variations in the manifestation of shame and pride, we computed the normalized distribution of psycho-social categories from the dialogues and examined their correlation with shame and pride compared to the control set. The search keywords used for building social emotions corpus (See Table A1) were removed from the LIWC dictionary to prevent overestimation of shame- (e.g., negative emotion) and pride-related categories (e.g., achievement).
3.2 Pretrained LLM: GPT-4 Chat
LLMs such as GPT-4 exhibit superlative English language understanding and are more reliable than human annotators on platforms such as MTurk Pan et al. (2023); Törnberg (2023). We used GPT-4 chat in two-shot setting ( See Tables A5 and A6 for prompts) to seek answers to the following questions:
  • who is $<$being shamed/feeling proud$>$ in the given movie discourse, and what is their gender?
  • What is the reason behind $<$the feeling of shame/pride$>$?
Since there are at least two characters in a discourse, the first question orients the LLM to focus on the person experiencing the social sanction or approval and then identify their gender. The output for the second question serves as the implicit social norm in the culture. Asking for "reason" leading to the expression of social emotion encourages LLMs to retrieve norm from provided context, mitigating potential Anglocentric tendencies.
Thematic Analysis of Social Norms
To capture overarching themes in reasons behind expressing shame and pride in Indian and American societies, we performed agglomerative clustering after embedding unique shame and pride-related social norms using SBERT embeddings Reimers and Gurevych (2019). Unlike movie dialogues, social norms extracted from GPT-4 chat are short phrases (See Table 4) and are devoid of culture-specific language style markers (i.e., Indian English vs American English) thus, no control for culture is needed.
Hollywood: LIWC Category (Top-5 words) $r$ Bollywood: LIWC Category (Top-5 words) $r$
Social References (you, your, we, he, you’re) 0.131 Drive (we, our, us, sir, work) 0.122
Family (son, dad, baby, mom, mama) 0.121 Achieve (work, better, win, best, try) 0.098
$1^{st}$ person sing. pronouns (i, i’m, me, my, i’ll) 0.095 Power (sir, own, respect, kill, power) 0.091
$2^{nd}$ person pronouns (you, your, you’re, you’ve, yourself) 0.091 Social References (you, your, he, we, our) 0.085
Male References (he, his, him, man, son) 0.084 Moral (wrong, duty, brave, arrogant, useless) 0.075
Drive (we, our, us, we’re, dad) 0.076 Reward (win, won, glory, success, successful) 0.071
Ethnicity (American, Irish, Chinese, German, Christian) 0.052 Political (nation, army, sultan, president, dynasty) 0.064
Reward (win, won, winner, successful, earned) 0.048 $1^{st}$ person pl. pronouns (we, our, us, we’ll, let’s) 0.060
Achieve (work, better, best, trying, try) 0.045 Family (son, papa, married, dad, uncle) 0.059
Power (sir, own, war, strong, mighty) 0.044 Ethnicity (Indian, Indians, British, Hindi, caste) 0.054
Table 3: Pearson r for LIWC categories significantly correlated ($p<0.05$) with pride for Hollywood and Bollywood after Benjamini-Hochberg p-value correction (See Appendix for confidence intervals and p values). Note the contrast, Achievement-related pride in Bollywood, Self-centered pride vs We-centered pride, Family in Hollywood vs Family in Bollywood See Tables A8 and A10 for all significantly correlated categories.
4 Results
4.1 How do expressions of shame/pride vary across cultures?
Shame is associated with negative emotions, power, and morality in both movie industries (See Table 2), however, significant cultural nuances exist in its manifestation. In Hollywood, shame is more I-focused with remorse (e.g., sadness, anxiety) and past-focused language, whereas in Bollywood, it is more You-focused with frequent social references and present-focused language. In Bollywood, psycho-social categories associated with shame reveal a larger social context (note the contrast in family i.e., nuclear family in Hollywood vs extended family in Bollywood) and social situations (i.e., anger, swear, conflict). Categories female and sexual are exclusively correlated with shame in Bollywood potentially indicating an honor system in collectivist societies Caffaro et al. (2014) whereas it is illness that is uniquely correlated with shame in Hollywood indicating social sanctions around incapability.
A similar dichotomy is observed for Pride related interactions (See Table 3). Pride in Hollywood interactions is more centered on family (note, family in Bollywood is male-centric and comprises extended family (such as uncle)) and self-centered (e.g., I am proud of __) whereas it is achievement/power focused and "We"-centered in Bollywood movies (with othering (us vs they)- See Appendix). Male references are more strongly associated with pride than females in Hollywood whereas the female category has an insignificant ($p>0.05$) correlation with pride in Bollywood. Ethnicity-related pride is expressed in both movie industries but political category is exclusively correlated to Bollywood. Pride is more positive (see POSEMO in Table 3) in Hollywood whereas in Bollywood, it reflects honor (e.g., protecting family’s pride, bringing pride to family).
Bollywood
eavesdropping on private conversation
expressing love for a man
Incestuous relationship
giving birth to a girl child
Poor academic performance
Hollywood
not living up to expectations
hiding/avoiding confrontation
not returning calls after intimacy
mistreatment of a woman
offering poor quality goods for sale
Table 4: A subset of reasons extracted from movie dialogues expressing shame. A total of 4604 unique reasons (Bollywood-3660, Hollywood-944) were extracted.
4.2 How do reasons behind shame/pride vary across cultures?
Cross-Cultural Norms Dataset
For the Bollywood set, GPT-4 chat predicted reasons (also implicit social norms) for 5321 (98.4%) shame-related dialogues out of 5409, and 2237 (74.6%) pride-related dialogues out of 2999. For the Hollywood set, GPT-4 chat predicted a reason for 1156 (94.6%) shame-related dialogues out of 1221 and 1731 (61.7%) pride-related dialogues out of 2805. Upon manual analysis, we found that pride is also used to express affection, specifically toward close family members such as sons - therefore, a lower number of norms is associated with pride compared to shame. GPT-4 chat predicted a total of 10,445 social norms (See Table A3 for norm distribution). Prompting "reasons" behind shame and pride allowed us to capture high specificity in cultural norms (See Tables 4 and 6 for examples).
4.2.1 Manual Evaluation
Two volunteers manually verified the predicted gender for "the person experiencing shame/pride" and the predicted "reason" in a randomly sampled set of $100$ dialogues each from Hollywood and Bollywood. The annotator who labeled the Bollywood set is an Indian, aware of social roles and expectations in Indian society. Likewise, the annotator for the Hollywood set is an American with a nuanced understanding of the social norms of the U.S.A.
For Bollywood, 8% of predicted gender and 11% of predicted reasons were incorrect where it was 5% and 2% respectively for Hollywood (See Table 5). There were 20 cases where gender (15) and/or reason (9) were ambiguous whereas, for Hollywood, there were 11 (gender=10, reason=1) such samples.
Gender Evaluation
Incorrect Ambiguous
Bollywood 8 15 (3 Female)
Hollywood 5 10 (3 Female)
Social Norms/Reason Evaluation
Incorrect Ambiguous
Bollywood 11 9
Hollywood 2 1
Table 5: Manual Evaluation of predicted gender and reasons for randomly sampled 100 samples each from Bollywood and Hollywood.
Bollywood
mastering a trick
fulfilling father’s dreams
provide care for old parents
fiancee’s physical appearance
his wealth
Hollywood
for being a hard worker regardless of the task
being a brilliant student
winning olympic gold
achievements and growth
coming out as queer
Table 6: A subset of reasons extracted from movie dialogues expressing pride. A total of 3163 unique reasons (Bollywood-1589, Hollywood-1574) were extracted.
4.2.2 Thematic Analysis
Twenty-five clusters for shame-related norms and 13 for pride-related norms were formed using agglomerative clustering ( See Tables A13 and A14 for clustering parameters). The clusters were manually assigned a theme (as depicted on the Y-axis of Fig. 2 and 3) after analyzing the ten closest samples based on cosine distance within the cluster (See Table A11 and A12 for examples). We computed the relative association for each cluster with Bollywood and Hollywood using eq. 1 and performed the Barnard-Exact Test Barnard (1947) with the Yates Correction Yates (1934) to test if the possibility of observing norms related to pre-assigned themes is statistically different across movie industries.
$\vec{\Delta}=\forall_{t_{i}\in themes}\frac{\mathcal{D}_{bolly_{t_{i}}}}{{\mathcal{D}_{bolly}}}-\frac{\mathcal{D}_{holly_{t_{i}}}}{\mathcal{D}_{holly}}$ (1)
Here, $\mathcal{D}_{industry_{t_{i}}}$ represents dialogues to $industry\in\{holly,bolly\}$ and ${t_{i}}\in themes$ (as depicted on Y-axis in Fig 2 and 3).
Figure 2: Relative association ($\Delta$) of Bollywood and Hollywood to themes obtained from agglomerative clustering performed on shame-related norms. * indicates insignificant difference i.e., $p>0.05$
Shame-related social norms
Themes such as lack of accountability and poverty are more common in Hollywood whereas norms related to sexual behavior and gender roles are more prevalent in Bollywood (See Fig. 2). As expected, collectivist factors such as non-conformity in gender roles, disrespect, and deviation from family norms are strongly associated with shame in Bollywood whereas individualistic attributes such as poverty, poor social etiquette and incompetence evoke shame in Hollywood.
Pride-related social norms
Duty, doing the "right" thing, and, achievement are associated with pride in Hollywood whereas Ethnolinguistic identity, family honor, and son’s success are associated with pride in Bollywood (See Fig. 3).
Figure 3: Relative association ($\Delta$) of Bollywood and Hollywood to themes obtained from agglomerative clustering performed on pride-related norms. * indicates insignificant difference i.e., $p>0.05$
4.2.3 Gender differences in Social Emotions
For the Bollywood set, GPT-4 predicted 1541 targets as female and 3949 as male. For the Hollywood set, GPT-4 predicted 482 targets as female and 1367 as male. Across all combinations (shame vs pride x Bollywood vs. Hollywood in Table A3), there are more male targets than females in line with the findings of Geena Davis Inclusion Quotient [^4] https://about.google/intl/ALL_us/main/gender-equality-films/.
Using predicted gender, we computed the gender-wise attribution to emotions "shame" and "pride" using eq. 2. A positive score indicates a higher association of gender groups with pride, whereas a negative score reflects a higher association with shame. A null score indicates no preference.
$\vec{\Delta}_{g}=\forall_{g\in\{male,female\}}\frac{\mathcal{D}_{pride_{g}}-\mathcal{D}_{shame_{g}}}{\mathcal{D}_{g}}$ (2)
Here, $\mathcal{D}_{pride_{g}}$ represents movie dialogues having pride-related word where the target of pride is $g\in\{male,female\}$, $\mathcal{D}_{shame_{g}}$ represents movie dialogues having shame-related word where the target of shame is $g\in\{male,female\}$ and $\mathcal{D}_{g}$ represents total movie dialogues grouped based on gender. As depicted in Fig. 4, Hollywood movies are pride-oriented, whereas Bollywood movies are shame-oriented. Females are attributed more shame, and the difference (male-female) in the expression of pride and shame is $0.16$ for Hollywood and $0.21$ for Bollywood.
Normative expectations evoking shame are similar for females across movie industries (See Fig. A1). Sexuality (e.g., promiscuity, immodesty) is the dominant theme for women in both movie industries whereas males are shamed for incompetency in both industries. Women across movie industries express more pride in family roles. Men-pride in Bollywood is centered on justice, winning and bravery whereas it is duty, self-identity and winning in Hollywood (See Fig. A2).
Figure 4: Relative association ($\Delta$) of emotions pride and shame gender-wise in Hollywood and Bollywood dialogues. A higher positive score indicates a stronger association of gender with pride.
5 Discussion
Detecting social emotions
We release a cross-cultural dataset of shame (6477 samples) and pride (3968 samples) related discourse with over 10K social norms. To the best of our knowledge, the only other corpus having samples related to social emotions is GoEmotions dataset Demszky et al. (2020) with 817 samples for embarrassment and 452 samples for pride. Detecting social emotions, and recognizing how appraisals and expressions of such emotions vary will improve human-computer communications, specially in domains of AI powered mental health therapy Kim et al. (2011), cross-cultural intent translation [^5] DARPA Computational Cultural Understanding: https://www.darpa.mil/program/computational-cultural-understanding.
Cross-cultural variations in expressions of social emotions
Shame is a highly undesirable self-focused emotion in the U.S. emphasizing incompetency and failures and is rarely used Cohen (2003); Boiger et al. (2013). Notably, we also observe its infrequent use in Hollywood movies and self-focus (reflecting internal shame) (See Table A2 and 2). In contrast, shame is considered interdependent in collectivist communities Wong and Tsai (2007). Our empirical analyses reveal You-focused (e.g. You, social references) or public nature of shame in Bollywood movies (See Table A7). Moreover, the contrasting tenses coupled with varying emotions, i.e., past + sadness in Hollywood vs. present + anger in Bollywood, reflect their varying goals, i.e., remorse for past failures/losses vs enforcing conformity Wong and Tsai (2007). Pride-related discourse in Hollywood is duty and achievement-focused, in line with prior findings underlining the increasing significance of "success" with growing individualism Cohen (2003), whereas in Bollywood, pride is centered around collective achievement (we, achievement, prosocial). It is also interesting to note linguistic markers such as determiner (used with objects/nouns) in Bollywood indicating more materialistic pride, compared to personal pronouns and social references in Hollywood reflecting more personal pride (See Tables A8, A10).
Cross-cultural variations in social sanctions and approvals
Eliciting reasons leading to social emotions shame and pride revealed socially sanctioned behaviors across cultures. We observed high specificity in extracted social norms revealing cultural subtleties (e.g., desire for a son, parents’ duty to marry off their daughter in Bollywood vs. honesty in business, returning calls after date night in Hollywood as in Table 4). We acknowledge that the excerpts explicitly expressing shame and pride from movie subtitles may have led to the over-representation of certain social situations (e.g., son’s achievement and daughter’s wedding in India vs duty and competence in America). A natural next step would be to capture situations with implicit shame and pride. The pipeline in Fig. 1 could be adapted to first detect self-conscious emotion in a given discourse and then identify the reason for advanced computational cultural understanding of the text. For this purpose, our dataset of shame/pride-related situations will be useful.
To tease apart ways in which cultures differ, we first considered mapping social norms to Schwartz’s Theory of Values Schwartz (2012). However, a social norm can have contrasting values depending on the culture. Consider "refusing to marry", an instance of non-conformity in Indian society, whereas it is an instance of self-direction in a Western context. Of course, society is evolving and we acknowledge that urbanized regions of India may also consider this an instance of self-direction. However, the diachronic shift in the underlying values of social norms is out of the scope of this paper. To effectively understand the cross-cultural variations in social norms, we performed hierarchical clustering and empirically picked the distance after manually analyzing the quality and granularity of clusters. The situations (i.e., poverty, inappropriate behavior, incompetence) connected to shame in the U.S. society vs situations(i.e., gender roles, disrespect, family norms) in Indian society as evident in Fig 2 align with known cultural tendencies of both nations. Likewise, pride is strongly associated with duty and achievement in Hollywood whereas with collective achievement (e.g., Ethnolinguistic identity) in Bollywood. While the clustering was performed on the reasons extracted from movie conversations, the cluster themes are very similar to patterns seen during LIWC analysis, supporting the use of social emotions for overcoming cultural bias in LLMs.
Women face more social sanctions than men
Shame, a self-conscious emotion expressing devaluation, is targeted towards women more than men, whereas pride, a positive valence emotion expressing value, is used for more men than women in movies. It is also surprising to see the similarity in normative expectations associated with genders male and female in Indian and American societies. Our analysis reveals that women are more frequent targets of social sanctions than men. It is thus important to characterize the source and acquired social norms before their use for achieving pluralistic values in LLMs otherwise we risk social bias.
6 Background
The vision of safe and accountable AI is centered on LLMs’ moral and value alignment. SOCIAL-CHEM-101 Forbes et al. (2020), SCRUPLES Lourie et al. (2021), SOCIAL BIAS FRAMES Sap et al. (2019), MORAL INTEGRITY CORPUS Ziems et al. (2022), and VALUE PRISM Sorensen et al. (2023) are a few datasets developed to teach socially aligned interactions to LLMs. The dominant approach for norm discovery involves prompting LLMs, sometimes coupled with a verification step such as an entailment test or underlying emotion (negative emotion $\rightarrow$ norm violation) Jiang et al. (2021); Fung et al. (2022); CH-Wang et al. (2023). However, most of the social situations in these datasets and the human annotators employed to label those situations reflect English beliefs and ethics. Additionally, language models prompted to identify "social norms" are known to reflect the values and beliefs of WEIRD nations (Western, Educated, Industrialized, Rich, and Democratic) Havaldar et al. (2023b); Atari et al. (2023).
7 Conclusion
We introduce an emotion-based approach to extract and analyze social norms across cultures. We create the largest known multi-cultural dataset of self-conscious emotions and relevant underlying norms. Our post hoc analysis of social norms demonstrates (a) cross-cultural linguistic differences in self-conscious emotions (shame and pride), (b) cultural dichotomy in normative expectations in India and the U.S.A., and (c) more "social sanctions" and fewer "social benefits" to the female gender in movies.
Future work can utilize our dataset to better align LLMs to various cultures. Additionally, our analysis of self-conscious emotions is the first of its kind; we hope future NLP researchers will build upon this work to investigate social norms in LLMs from a multicultural, emotion-based perspective.
Social Impact and Ethics Statement
Social norms discovery is a crucial component in programs [^6] ALIGN-https://www.alignplatform.org/learning-collaborative designed for social and behavioral change to promote equity, social justice, and well-being Mauduy et al. (2022); Bonan et al. (2020). Further work explores style as a product of norm differences Havaldar et al. (2023a). Social psychology investigates social norms (descriptive vs injunctive) to design experiments for understanding behaviors such as self-regulation, persuasion Cialdini et al. (1990) and decision-making Gavrilets (2020); Bhanot (2021) to promote collective-level change in societies.
Relatedly, Kimbrough and Vostroknutov (2023) showed people’s tendency to choose self-serving social norms using a dictator-recipient setup, emphasizing the need for dedicated research efforts to understand morality and belief distortion in different contexts. The norms and cultural preferences learned from movies that often showcase stereotypical behaviors of society may induce pluralistic ignorance and, more importantly, lead to discrimination and biases in LLMs when used for training. We hope that this paper will encourage scrutiny of source corpora and derived norms before their use for fine-tuning LLMs.
Limitations
Social norms mutate as society evolves. We acknowledge that our dataset of movies (released post-1990) may reflect social norms that are less characteristic of contemporary society. Moreover, countries like India and America contain a mix of cultures. The captured norms may not reflect the cultural variations, for example, between regions (e.g., East Coast vs West Coast in the U.S.A or North India vs South India). Movies also exaggeratedly depict the world around us (e.g., weddings, criminal activities, sexual abuse, etc.), and we caution against stereotyping cultures based on movie-based norms.
The dominant language in Bollywood movies is Hindi and our analysis is based on their English translations which may not always be accurate, especially when the discourse is about concepts native to a culture. Relatedly, the LIWC may not have high coverage for such concepts. We did not compare the movie genre and acknowledge that situational/unrealistic norms (e.g., science fiction, comedy, etc.) could exist. Regardless, social norms associated with shame and pride are still relatable and reflect the target audience’s beliefs. It is also worth noting that gender norms in movies may not fully reflect the norms of today’s society. This study is conducted over a thirty-year long period and thus, the aggregated norms may not reflect the current trends.
The prompts ("You study Indian society" vs "You study Western society" described in Tables A5 and A6 ) used to elicit norms may have induced unwanted bias Cheng et al. (2023); Lucy and Bamman (2021) and it is worth investigating the variations if any, in extracted norms due to different prompt designs.
Acknowledgements
We acknowledge the use of ChatGPT, an AI language model developed by OpenAI, for its assistance in writing and refining this manuscript.
Appendix A Appendix
A.1 Search Keywords
Search keywords defined in Table A1 were used to extract shame- and pride-related dialogues from Bollywood and Hollywood movie subtitles.
Lemma Word forms
shame shame, shamed, shameful, ashamed
pride proud, prouder, proudly, pride
Table A1: List of search keywords for extracting dialogues.
Movies Total Tokens #shame #pride #control
Bollywood 2738 22.62M 5409 2999 8303
Hollywood 2697 20.78M 1221 2805 4385
5435 43.3M 6583 5804 12688
Table A2: Data Distribution. #shame indicates the number of dialogues with the word "shame" or its derivative form (e.g., ashamed, shameless). Similarly, #pride indicates the number of dialogues with the word pride or its derivative form (e.g., proud). Control is the group of dialogues without words shame, pride, and their derivatives.
A.2 Examining Implicit shame:
A.3 Norms Distributions
Over 10k reasons (or norms) were extracted using GPT-4. The distribution of norms and gender of targets is provided in Table A3.
Bollywood Hollywood
Gender shame pride shame pride
male 2690 1259 591 776
female 1215 326 246 236
Not clear 1416 652 319 719
total 5321 2237 1156 1731
Table A3: Norm and Gender distribution (with duplicates) for dialogues for which GPT-4 predicted male or female. The duplicate norms are not removed as their frequency reflect their prevalence and is useful for estimating gender association.
A.4 Annotation
The annotation guidelines to verify the gender predicted by GPT-4 and the correctness of the reason are provided in Table A4. The annotators for Bollywood set and Hollywood set were Indian and American respectively. Both annotators were female, proficient in English language, and well-versed with social norms. During annotation, if the gender or the reason is unclear, the annotators were asked to label "not explicit". We only considered the cases where the gender was predicted to be either male or female. The task is objective and inter-annotator agreement was not computed. The annotators volunteered for the task and were not provided monetary compensation.
Guidelines for Manual Evaluation
1. Read the conversation and identify the person feeling ashamed (or being shamed) or proud.
2. Identify the gender. Check gender markers such as Mr/Mrs., s/he, him/her, etc. If the name is provided in the conversation, check if the name is likely to be a male name or female. If not clear, mark "not explicit".
3. Read the reason behind shame/pride. Compare with conversation and determine if the provided reason is the cause for shame/pride.
Table A4: Guidelines for Annotation
A.5 Prompts
The prompts designed to extract the subject experiencing self-cosncious emotions, their gender and the reason behind the emotions are provided in Tables A5 and A6.
Prompts for norm discovery in Bollywood movie dialogues
"You are an experienced social scientist and you study Indian society. Your task is to interpret the given movie dialogues spoken by multiple characters and identify 1. Who is feeling proud? Provide the gender. If agent or gender is not explicit, use the pronouns, name, and salutations to guess agent and gender. 2. Identify the reason behind feeling proud. If not explicit, return NA. The answer should be short and in below CSV format. $<$who_proud, gender, reason$>$
Input: Your mother had already given up on me because I was useless hopeless But you were the apple of her eye My child will make me proud look after me He’ll support in my old-age But this kid left her all alone.
Output: mother, male, provide care for old parents Input: I’ve heard there’s a promising young student in your school What’s his name? He’s made us proud in long jump, we are here to felicitate him Call him Show yourself, Raju Tempre Output: authority, NA, Sports achievement"
"You are an experienced social scientist and you study Indian society. Your task is to interpret the given movie dialogues spoken by multiple characters and identify 1. Who is being shamed? Predict the gender. If gender is not explicit, use the pronouns, name, and salutations to guess gender. 2. Identify the primary reason for shaming The answer should be short and in CSV format. $<$ shamed, gender, reason$>$ Input: And should we bow before others begging....them to marry our daughters? This shall not happen. Neither will the girls be alive here nor shall....we be ashamed of ourselves. You cannot kill the life which God has given. I won’t let you commit the sin. Output: girl’s parent, NA, not able to marry off their daughters Input: Black marketers are now in the open. And the thieves too Politics is in a great mess Shame on this system. There’s no democracy Get rid of these politicians The gong has struck..’’Our hearts are swaying to it’s beats’’ Output: System, NA, poor law and regulations"
Table A5: Bollywood: Prompts for norm discovery using GPT-4 Chat. The temperature was set to 00 to minimize randomness.
Prompts for norm discovery in Hollywood movie dialogues
"You are an experienced social scientist and you study Western society. Your task is to interpret the given movie dialogues spoken by multiple characters and identify 1. Who is feeling proud? Provide the gender. If agent or gender is not explicit, use the pronouns, name, and salutations to guess agent and gender. 2. Identify the reason behind feeling proud? If not explicit, return NA. The answer should be short and in below CSV format. $<$who_proud, gender, reason$>$ Input: I want to go to Worlds and win gold. I want to go to the 88 Olympics in Seoul and win gold. Good! I’m proud of you. Are you getting the support that you need? What do you mean sir? Output: Sir, male, winning olympic gold Input: Yes. Yes, I did. I promise, this time I really got the promotion. - I’m proud of you, son. - Thank you, sir. Excuse me. Hi, sweetheart. Output: father, male, for getting the promotion"
"You are an experienced social scientist and you study western societies. Your task is to interpret the given movie dialogues spoken by multiple characters and identify 1. Who is being shamed? Predict the gender. If gender is not explicit, use the pronouns, name, and salutations to guess gender. 2. Identify the primary reason for shaming. The answer should be short and in CSV format. $<$ shamed, gender, reason, prevalence $>$ Input: You still owe me 100. Remember? You stiffed Donny for 100 bucks? Cheapskate. Shame on you. Pay this man his C-note. Now I know why they call you the Snake. Output: NA, male, not returning borrowed money Input: You prey on your own people. You steal from your own people. Have you no shame!? - Huh? - Well, we’re still here. Man: Mr. Markopolos, it’s all yours. Output: Snake, male, stealing and preying on people"
Table A6: Hollywood: Prompts for norm discovery using GPT-4 Chat. The temperature was set to 00 to minimize randomness.
A.6 LIWC Correlation Results
Tables A7, A8, A9 and A10 contain the significantly correlated ($p<0.05$) LIWC categories, the most frequent five words for each category, pearson r and 95% confidence interval.
LIWC Categories Top-5 words r p-value 95% CI
Negative emotions (EMO_NEG) (bad, mad, scared, worry, fear) 0.330 0.000 [0.315, 0.344]
Negative tone (TONE_NEG) (lost, kill, wrong, bad, hit) 0.249 0.000 [0.233, 0.265]
POWER (sir, respect, own, kill, poor) 0.198 0.000 [0.182, 0.214]
EMOTION (love, good, bad, happy, crazy) 0.168 0.000 [0.152, 0.185]
YOU (you, your, you’re, yourself, you’ve) 0.161 0.000 [0.145, 0.178]
Social references (SOCREFS) (you, your, he, her, him) 0.148 0.000 [0.132, 0.165]
FEELING (feel, touch, feeling, felt, hard) 0.133 0.000 [0.117, 0.150]
DRIVES (we, our, us, sir, married) 0.115 0.000 [0.098, 0.132]
SOCIAL (you, your, he, her, him) 0.110 0.000 [0.094, 0.127]
MORAL (wrong, innocent, duty, decent, excuse) 0.102 0.000 [0.085, 0.118]
AFFECT (love, good, keep, respect, well) 0.078 0.000 [0.062, 0.095]
NEGATE (not, don’t, no, aren’t, won’t) 0.065 0.000 [0.049, 0.082]
FEMALE (her, she, girl, she’s, mom) 0.060 0.000 [0.043, 0.077]
Personal Pronouns (PPRON) (you, i, me, your, my) 0.058 0.000 [0.041, 0.075]
FAMILY (son, married, uncle, dad, mom) 0.055 0.000 [0.039, 0.072]
Preposition (PREP) (to, of, in, for, on) 0.053 0.000 [0.037, 0.070]
SEXUAL (chaste, lust, sex, sexy, pimp) 0.051 0.000 [0.034, 0.067]
PRONOUN (you, i, me, your, my) 0.040 0.000 [0.023, 0.057]
AUXVERB (is, are, have, be, don’t) 0.029 0.001 [0.013, 0.046]
CONFLICT (kill, killed, accusing, killing, cruel) 0.024 0.007 [0.007, 0.041]
SWEAR (hell, bloody, idiot, damn, ass) 0.024 0.007 [0.007, 0.041]
Anger (EMO_ANGER) (mad, angry, hate, cruel, argue) 0.022 0.015 [0.005, 0.039]
FOCUSPRESENT (is, are, don’t, i’m, aren’t) 0.021 0.018 [0.005, 0.038]
Table A7: Psychosocial categories significantly correlated (p<0.05) with shame in Bollywood dialogues. p-values were corrected using Benjamini-Hochberg correction. The categories are arranged in decreasing order of correlation.
LIWC Categories Top-5 words r p-value 95% CI
DRIVES (we, our, us, sir, work) 0.122 0.000 [0.104,0.140]
Determiner (DET) (the, a, my, your, that) 0.101 0.000 [0.082,0.119]
ACHIEVE (work, better, win, best, try) 0.098 0.000 [0.080,0.116]
POWER (sir, own, respect, kill, power) 0.091 0.000 [0.072,0.109]
Social References (SOCREFS) (you, your, he, we, our) 0.085 0.000 [0.067,0.103]
Preposition (PREP) (of, to, in, for, with) 0.084 0.000 [0.066,0.103]
MORAL (wrong, duty, brave, arrogant, useless) 0.075 0.000 [0.057,0.093]
Conjunction (CONJ) (and, but, so, if, as) 0.075 0.000 [0.056,0.093]
REWARD (win, won, glory, success, successful) 0.071 0.000 [0.052,0.089]
Positive tone (TONE_POS) (love, good, thank, well, great) 0.064 0.000 [0.046,0.082]
POLITIC (nation, army, sultan, president, dynasty) 0.064 0.000 [0.045,0.082]
WE (we, our, us, we’ll, let’s) 0.060 0.000 [0.042,0.079]
SOCIAL (you, your, he, we, our) 0.060 0.000 [0.041,0.078]
FAMILY (son, papa, married, dad, uncle) 0.059 0.000 [0.041,0.078]
AFFILIATION (we, our, us, dear, we’ll) 0.059 0.000 [0.041,0.077]
FEELING (feel, feeling, hard, felt, sense) 0.055 0.000 [0.036,0.073]
ETHNICITY (indian, indians, british, hindi, caste) 0.054 0.000 [0.036,0.072]
MALE (he, his, him, son, sir) 0.054 0.000 [0.036,0.072]
CULTURE (indian, nation, army, car, indians) 0.044 0.000 [0.026,0.062]
AFFECT (love, good, thank, well, great) 0.040 0.000 [0.021,0.058]
ARTICLE (the, a, an, that) 0.039 0.000 [0.021,0.058]
Personal Pronouns (PPRON) (you, i, my, your, me) 0.036 0.000 [0.017,0.054]
PROSOCIAL (thank, please, sorry, respect, gift) 0.032 0.001 [0.014,0.051]
FUNCTION (you, the, i, of, to) 0.032 0.001 [0.013,0.050]
YOU (you, your, you’re, you’ve, you’ll) 0.030 0.003 [0.011,0.048]
THEY (they, their, them, they’re, they’ll) 0.029 0.003 [0.011,0.048]
CERTITUDE (really, real, surely, proved, actually) 0.021 0.041 [0.002,0.039]
Positive Emotion (EMO_POS) (love, good, happy, happiness, smile) 0.020 0.049 [0.001,0.038]
Table A8: Psychosocial categories significantly correlated (p<0.05) with pride in Bollywood dialogues. p-values were corrected using Benjamini-Hochberg correction. The categories are arranged in decreasing order of correlation.
LIWC Categories Top-5 words r p 95% CI
Negative emotions (EMO_NEG) (sick, pain, fear, bad, afraid) 0.425 0.000 [0.403, 0.446]
Negative tone (TONE_NEG) (lost, wrong, sick, pain, poor) 0.355 0.000 [0.331, 0.377]
EMOTION (good, love, sick, pain, bad) 0.290 0.000 [0.266, 0.314]
POWER (own, sir, poor, killed, war) 0.263 0.000 [0.238, 0.287]
AFFECT (well, good, love, help, damn) 0.168 0.000 [0.142, 0.193]
DRIVES (we, us, our, work, we’re) 0.152 0.000 [0.127, 0.178]
FUNCTION (you, i, the, to, of) 0.131 0.000 [0.105, 0.157]
Personal Pronouns (PPRON) (you, i, me, i’m, my) 0.111 0.000 [0.085, 0.137]
MORAL (wrong, excuse, decent, honest, duty) 0.110 0.000 [0.084, 0.136]
I (i, me, i’m, my, i’ll) 0.102 0.000 [0.076, 0.128]
Sadness (EMO_SAD) (crying, cry, sob, lonely, sad) 0.096 0.000 [0.070, 0.122]
NEGATE (no, not, don’t, nothing, never) 0.090 0.000 [0.063, 0.116]
Preposition (PREP) (to, of, in, for, on) 0.089 0.000 [0.062, 0.115]
PRONOUN (you, i, that, it, me) 0.088 0.000 [0.062, 0.114]
YOU (you, your, you’re, yourself, you’ve) 0.080 0.000 [0.054, 0.106]
Auxiliary verb (AUXVERB) (be, i’m, is, was, have) 0.077 0.000 [0.051, 0.103]
FOCUSPAST (was, did, were, been, didn’t) 0.069 0.000 [0.043, 0.096]
SOCIAL (you, your, we, he, you’re) 0.060 0.000 [0.033, 0.086]
Conjunction (CONJ) (and, so, but, if, when) 0.056 0.000 [0.030, 0.082]
LINGUISTIC (you, i, the, to, of) 0.056 0.000 [0.030, 0.082]
ALLNONE (no, all, nothing, never, yes) 0.055 0.000 [0.029, 0.081]
FAMILY (son, dad, baby, mom, mama) 0.054 0.000 [0.028, 0.081]
Social references (SOCREFS) (you, your, we, he, you’re) 0.050 0.000 [0.024, 0.077]
Anxiety (EMO_ANX) (fear, afraid, worry, terrified, scared) 0.042 0.004 [0.015, 0.068]
ILLNESS (sick, pain, pains, flu, sickly) 0.038 0.010 [0.012, 0.064]
FEELING (feel, felt, pain, feeling, hard) 0.035 0.020 [0.008, 0.061]
Anger (EMO_ANGER) (hate, hated, mad, angry, hates) 0.033 0.026 [0.007, 0.059]
DIFFER (not, but, if, didn’t, or) 0.033 0.027 [0.007, 0.059]
Discrepancy (DISCREP) (should, can, would, can’t, want) 0.032 0.029 [0.006, 0.059]
COGNITION (no, not, all, know, but) 0.032 0.033 [0.005, 0.058]
Table A9: Psychosocial categories significantly correlated (p<0.05) with shame in Hollywood dialogues. p-values were corrected using Benjamini-Hochberg correction. The categories are arranged in decreasing order of correlation.
LIWC Categories Top-5 words r p-value 95% CI
Personal Pronouns (PPRON) (you, i, i’m, me, my) 0.135 0.000 [0.113,0.158]
Social References (SOCREFS) (you, your, we, he, you’re) 0.131 0.000 [0.108,0.154]
FAMILY (son, dad, baby, mom, mama) 0.121 0.000 [0.099,0.144]
Conjunction (CONJ) (and, so, but, as, if) 0.114 0.000 [0.091,0.137]
SOCIAL (you, your, we, he, you’re) 0.102 0.000 [0.079,0.125]
FUNCTION (you, i, the, of, to) 0.100 0.000 [0.077,0.123]
I (i, i’m, me, my, i’ll) 0.095 0.000 [0.072,0.118]
YOU (you, your, you’re, you’ve, yourself) 0.091 0.000 [0.068,0.114]
MALE (he, his, him, man, son) 0.084 0.000 [0.061,0.107]
DRIVES (we, our, us, we’re, dad) 0.076 0.000 [0.053,0.099]
Positive tone (TONE_POS) (good, well, thank, great, love) 0.072 0.000 [0.049,0.095]
PRONOUN (you, i, i’m, that, it) 0.072 0.000 [0.048,0.095]
Auxiliary Verb (AUXVERB) (i’m, be, is, was, have) 0.071 0.000 [0.048,0.094]
Positive Emotion (EMO_POS) (good, love, happy, hope, wonderful) 0.070 0.000 [0.047,0.093]
Preposition (PREP) (of, to, in, for, on) 0.070 0.000 [0.047,0.093]
AFFILIATION (we, our, us, we’re, dad) 0.054 0.000 [0.031,0.078]
EMOTION (good, love, happy, hope, bad) 0.052 0.000 [0.029,0.075]
ETHNICITY (american, irish, chinese, german, christian) 0.052 0.000 [0.029,0.075]
REWARD (win, won, winner, successful, earned) 0.048 0.000 [0.025,0.071]
ACHIEVE (work, better, best, trying, try) 0.045 0.000 [0.022,0.068]
POWER (sir, own, war, strong, mighty) 0.044 0.001 [0.021,0.067]
AFFECT (good, well, thank, great, love) 0.041 0.001 [0.018,0.064]
MORAL (wrong, excuse, hero, brave, dignity) 0.040 0.002 [0.017,0.064]
FEMALE (her, she, she’s, girl, ladies) 0.034 0.009 [0.011,0.057]
CULTURE (american, car, president, nation, mayor) 0.033 0.012 [0.009,0.056]
FOCUSPAST (was, did, been, were, had) 0.032 0.013 [0.009,0.055]
SHEHE (he, his, him, her, she) 0.031 0.016 [0.008,0.054]
WORK (work, job, school, deal, company) 0.028 0.033 [0.005,0.051]
Table A10: Psychosocial categories significantly correlated (p<0.05) with pride in Hollywood dialogues with 95% confidence intervals. p-values were corrected using Benjamini-Hochberg correction. The categories are arranged in decreasing order of correlation.
A.7 Clustering Results
Tables A13 and A14 contain the manually annotated Cluster Theme, the total number of samples in each cluster, and Bollywood vs Hollywood distribution. The distance was set to $5$ and the duplicates were removed. The theme and top three examples demonstrating its meaning are provided in Tables A11 and A12.
Theme Examples from cluster
Stealing shamelessness and taking money, lying and misusing money, Stealing and bribery
Poverty not having money, being poor, living in poor conditions
Incompetence not fulfilling responsibilities, not meeting expectations or making a mistake, Failure or perceived incompetence
Promiscuity being promiscuous, Having a love affair, shameless behavior and expressing love
Inappropriate behavior not living up to someone’s expectations, being disloyal, Being rude and ungrateful
Disobedience not upholding cultural values, disobedience or lack of respect, being disloyal
Parent-related not standing up for his mother, not taking care of his father, not taking responsibility for his son
Marriage-related refusing to marry, being forced into marriage, not accepting the proposed marriage
Immodesty Wearing inappropriate clothes, inappropriate behavior nudity in public, behaving indecently in public
Gender Roles not behaving as per her husband’s expectations, behaving inappropriately in front of her daughter, going against her husband
Sexual Harrassment assaulting a girl, physical assault on women, sexual assault
Cowardice lack of courage, cowardice and inability to stand up for oneself, lack of pride and integrity
Alcholism excessive drinking, inappropriate behavior due to alcohol, drinking and irresponsible behavior
Non-conformity not conforming to gender norms, breaking gender norms, behaving inappropriately according to societal norms
Illegal activities engaging in criminal activities, Being a criminal, committing illegal acts
Betrayal Betrayal or dishonesty, betrayal and infidelity, Deception/Betrayal
Lack of accountability not taking responsibility and blaming others, wrongdoing without remorse, not acknowledging wrongdoing
Disrespect disrespecting others, disrespecting an authority figure, Disrespectful and inappropriate behavior
Harm causing harm to others, committing harmful deeds, causing trouble and endangering others
Lying/Deception lying and hiding information, Lying or deceit, Deception/Not being truthful
Inappropriate sexual behavior inappropriate advances and comments, inappropriate behavior towards a young girl, inappropriate language and behavior
Social Etiquette Being humiliated in public, being disrespected and belittled, being mocked and treated shamefully
Privacy-related invading personal space, invading someone’s privacy, intrusion of personal space, trespassing/invading personal space
Accusation accused of wrongdoing, being accused of something, being accused of infidelity
Family norms Disrespecting family, causing harm and shame to family, causing difficulty and shame to family
Table A11: Themes and top-3 examples for clusters obtained after agglomerative clustering of shame-related dialogues from Bollywood and Hollywood.
Theme Examples from cluster
Achievement being a brilliant student, working hard and achieving a spectacular result, professional achievement
Family Roles raising their child, having an unselfish mother, having a simple and good daughter
Self-identity being a man, being a woman, being gay
Duty fulfilling duty, Accomplishing something great, saving the world
Bravery & Nation bravery and sacrifice, bravery and service to nation, bravery and selflessness
Doing the "right" thing for doing the right thing, for being heroes, for having courage
Son’s success son’s success and progress, son’s hard work and achievement, son’s determination
Family Honor maintaining dignity and pride, taking care of family pride, saving family’s pride
Justice standing up for justice, fighting for justice, bringing justice
Marriage-related marrying with pride and respect, sister’s marriage, daughter’s marriage into a reputed family
Appearance & Assets his manhood, his wealth and beautiful wife, his knowledge and power
Resilience resilience and determination, overcoming struggles, enduring hardships without complaint
Ethnolinguistic Identity being Indian, defending the pride of Rajputs, being a Maharashtrian
Table A12: Themes and top-3 examples for clusters obtained after agglomerative clustering of pride-related dialogues from Bollywood and Hollywood.
Theme (Shame-related norms) Total Samples Hollywood Bollywood
Accountability 802 185 617
Accusation 70 0 70
Alcoholism 98 18 80
Betrayal 262 26 236
Cowardice 209 51 158
Disobedience 216 49 167
Disrespect 246 26 220
Family norms 203 24 179
Gender roles 221 17 204
Harm 225 64 161
Illegal activities 253 27 226
Immodesty 311 58 253
Incompetence 342 83 259
Lying/Deception 222 53 169
Marriage-related 179 25 154
Non-conformity 208 28 180
Parent-related 263 34 229
Poverty 414 108 306
Privacy-related 102 14 88
Promiscuity 595 89 506
Sexual behavior 407 33 374
Sexual Harassment 232 24 208
Social Etiquette 417 97 320
Stealing 351 49 302
Total 6848 1182 5666
Table A13: Distribution of reasons (shame) across manually labeled clusters. A total of twenty-six clusters were generated with distance=5. Duplicates were removed for clustering. One cluster with generic reasons ( such as phrases "lack of shame") was removed and two clusters with similar reasons (related to accountability) were merged. Finally, 24 clusters were considered. The total is slightly more than numbers in Table A3 since a norm could be mapped to multiple clusters.
Theme (Pride-related norms) Total Samples Hollywood Bollywood
Achievement 332 174 158
Bravery 248 69 179
Doing the "right" thing 82 82 0
Duty 792 449 343
Ethnolinguistic Identity 148 7 141
Family Honor 263 77 186
Family Roles 356 156 200
Justice 255 109 146
Nation 216 57 159
Physical Appearance 106 49 57
Resilience 180 68 112
Self-identity 273 157 116
Son’s Achievements 315 87 228
Wedding 120 23 97
Winning 411 205 206
Total 4097 1769 2328
Table A14: Distribution of reasons (pride) across manually labeled clusters. A total of fifteen clusters were generated with distance=5. Duplicates were removed for clustering. ’Achievement’ and ’Winning’ were merged to form ’Achievement’ and, ’Bravery’ and ’Nation’ were merged to form ’Bravery & Nation’. Finally, 13 clusters were considered for analysis.
(a) Shame-related norms in Bollywood
(b) Shame-related norms in Hollywood
Figure A1: Gender-wise differences in themes associated with normative expectations from shame-related discourse in Bollywood and Hollywood. $\Delta(Female,Male)$ is the difference between normalized dialogues attributed to females under a theme and to males under the same theme (as in eq. 1 but for dimension gender). * indicates insignificant difference i.e., $p>0.05$
(a) Shame-related norms in Bollywood
(b) Shame-related norms in Hollywood
Figure A2: Gender-wise differences in themes associated with normative expectations from pride-related discourse in Bollywood and Hollywood. $\Delta(Female,Male)$ is the difference between normalized dialogues attributed to females under a theme and to males under the same theme (as in eq. 1 but for dimension gender). * indicates insignificant difference i.e., $p>0.05$