### v4

# Social Norms in Cinema: A Cross-Cultural Analysis of Shame, Pride and Prejudice

Abstract
Shame and pride are social emotions expressed across cultures to motivate and regulate people’s thoughts, feelings, and behaviors. In this paper, we introduce the first cross-cultural dataset of over $10k$ shame/pride-related expressions, with underlying social expectations from  5.4K Bollywood and Hollywood movies. We examine how and why shame and pride are expressed across cultures using a blend of psychology-informed language analysis combined with large language models. We find significant cross-cultural differences in shame and pride expression aligning with known cultural tendencies of the USA and India -- e.g., in Hollywood, shame-expressions predominantly discuss self whereas Bollywood discusses shame toward others. Pride in Hollywood is individualistic with more self-referential singular pronouns such as I and my whereas in Bollywood, pride is collective with higher use of self-referential plural pronouns such as we and our. Lastly, women are more sanctioned across cultures and for violating similar social expectations e.g. promiscuity.
Social Norms in Cinema: A Cross-Cultural Analysis of Shame, Pride and Prejudice
1 Introduction
Shame and pride are social emotions expressed across cultures to motivate and regulate people’s thoughts, feelings, behaviors, achievements, and social adaptations Fischer and Tangney (1995); Goetz and Keltner (2007); Fessler (2007); Schaumberg and Skowronek (2022). The expressions and appraisals of shame and pride vary across cultures and is a research question of great significance to understand society and self-regulation Wong and Tsai (2007); Lewis et al. (2010); Wegge et al. (2022). The overarching question we ask in this paper is, how and why shame and pride are expressed across cultures. Knowing cross-cultural differences in shame/pride-related expressions will help us measure psychological constructs such as shame proneness Tangney et al. (1992); Cohen et al. (2011) and self-esteem Pyszczynski et al. (2004) reliably from language -- enabling effective human-computer interactions in domains such as AI-driven psychotherapy Resendiz and Klinger (2023); Demszky et al. (2020). Examining behaviors evoking shame and pride will reveal culture-specific societal expectations and attitudes (See Table 1) -- paramount to building culturally competent AI models for diverse users Talat et al. (2021); Atari et al. (2023).
Look, Alok is in love with me and I love him too.
Shameless one! Have you lost all your shame?
What is there to be ashamed in this? Don’t sisters love their brothers in this vicinity?
$\Rightarrow$ Incestuous relationship evoke shame.
Sister-in-law! Congrats, sister-in-law! Big brother has started working!
Really?
Now you will have a place of pride in this family!
Yes, please!
$\Rightarrow$ Employed husband evokes pride
Table 1: Excerpts of dialogues expressing explicit social emotions shame and pride indicating actions and social approval.
For our study, we select two top-grossing movie industries, i.e., Hollywood and Bollywood  [^1] https://nairametrics.com/2024/08/14/top10-fastest-growing-film-industries-in-the-world/. Hollywood primarily depicts social situations from the USA, an individualist society that values competency and autonomy Triandis (1989, 1988). Bollywood depicts India, a collectivist society where one’s sense of self is interwoven with community beliefs. The cultural dichotomy between India and the USA (i.e., collectivism vs. individualism) thus presents a rich ground for understanding variations in their beliefs and values. Our approach blends psychology-informed language analysis with state-of-the-art Large Language Model (LLM) to delineate interpretable psychosocial constructs such as self-focus, morality, etc. in shame/pride-related expressions and extract implicit social expectations behind them.
This paper makes the following research contributions: (a) We build a cross-cultural dataset of over 10k shame/pride-related dialogues with underlying social expectations  [^2] Link will be provided upon acceptance, (b) We demonstrate significant cross-cultural variations in how and why shame and pride are expressed in movies, and (c) We illustrate that women are more subjected to social sanctions than men across cultures and for violating similar social expectations.
To the best of our knowledge, this is the first empirical study analyzing the psycho-social constructs underlying shame/pride-related expressions across cultures and the social expectations behind them.
2 Movies Subtitles Corpora
Movies provide culture-specific life-like depictions of social situations Adkins and Castle (2014); Kubrak (2020). The natural conversation style between characters in movies can reveal the social power dynamics (e.g., boss-employee, father-daughter) and gender roles. Another potential alternative is social media posts however, they are (a) skewed to social situations specific to young demographic groups from Western countries, and (b) are second-hand reports that do not reveal fully the social dynamics of how shame/pride are expressed in natural conversations, for instance, between a boss and an employee or among family members at a Christmas table.
We collected English subtitles for 5,435 Hollywood and Bollywood movies that were released post-1990 by auto-crawling websites that host or link movie subtitles (See Table 2 for data distribution). The subtitles are professionally done translations to preserve the emotions and intended message. Bollywood movie subtitles were collected from www.Bollynook.com and Hollywood movie subtitles were from the publicly available Kaggle movie subtitles dataset [^3] https://www.kaggle.com/datasets/adiamaan/movie-subtitle-dataset. The year of release for movies was verified by either parsing subtitle file names having a release year or checking Wikipedia entries. The year mapping was performed to ensure a similar period for collected movies.
Hollywood Bollywood
#Movies 2697 2738
#Tokens 20.78M 22.62M
#Shame 1221 5409
#Pride 2805 2999
#Control 4385 8303
Table 2: Data Distribution. #shame indicates the number of dialogues with the word "shame" or its derivative form (e.g., ashamed, shameless - See Table A1). Similarly, #pride indicates the number of dialogues with the word pride or its derivative form (e.g., proud). Control is the group of dialogues without words shame, pride, and their derivatives.
2.1 Extracting Shame and Pride related Expressions
The expressions of shame and pride could be explicit (See Table 1) or implicit (e.g. characters’ lowered shoulders, avoiding eye contact, etc.). We extracted explicit mentions of shame and pride to learn about occurrences of shame and pride as determined by the characters. We adopted a keyword search-based approach (See Appendix A1) to identify dialogues along with the previous and the next two lines for situational context (See Table 1). In the case of short dialogues, which could be the case due to monosyllabic responses in spoken conversations, an extra previous and next line was appended to context. One author manually checked the entire dataset of shame and pride-related dialogues to filter out dialogues with phrases such as "what a shame", "it’s a shame", "proudly presents".
Detecting implicit shame/pride-related expressions in textual discourse is an innately complex task. Computational systems for detecting social emotions such as shame and guilt underperform significantly compared to basic emotions such as joy Demszky et al. (2020); Resendiz and Klinger (2023) due to internal self-evaluation associated with social emotions. We tested the GPT-4 language model Achiam et al. (2023) to detect implicit expressions of shame in given movie dialogues without explicit mentions of shame and pride. We recruited two human annotators to determine the agreement with GPT-4 generated labels (See Appendix A.2 for details). On a randomly sampled 100 situations each from H/Bollywood, the model returned 20 samples to be expressing shame in Bollywood and 5 samples in Hollywood. We found poor agreement between human annotators (4 out of 20 samples in Bollywood and 1 out of 5 samples in Hollywood where both annotators agreed with GPT-4). Human annotators across cultures tend to cross-label shame and guilt Troiano et al. (2019) and follow-up discussions with annotators revealed similar tendencies. One such example from Hollywood is provided below:
‘‘about feeding the poor, I’ve never done any of that. God tells us to love everybody. I’ve hated people… my family, my family, ’’
Here, the speaker is probably expressing guilt, shame, or both. Given subjective beliefs and poor agreement with generated outputs, we adopted a high precision-low recall strategy and limited our further analysis to explicit mentions of shame and pride.
For our analysis, we created four sets of dialogues: (a) shame-related dialogues in Bollywood, (b) shame-related dialogues in Hollywood, (c) pride-related dialogues in Bollywood and (d) pride-related dialogues in Hollywood. Cross-cultural differences also exist in speaking styles across cultures and we thus formed a control set of dialogues unrelated to shame and pride for both movie industries to remove the variations in language markers owing to culture-specific speaking styles (See Table 2 for dialogue distribution).
Figure 1: An overview of our approach comprising two key steps (a) Vocabulary approach and (b) Prompting a pre-trained LLM.
Hollywood: LIWC Category (Top-5 words) $r$ Bollywood: LIWC Category (Top-5 words) $r$
Moral (wrong, excuse, decent, honest, duty) 0.110 $2^{nd}$ person pronouns (you, your, you’re, yourself, you’ve) 0.161
$1^{st}$ person sing. pronouns (i, me, i’m, my, i’ll) 0.102 Social References (you, your, he, her, him ) 0.148
Sadness (crying, cry, sob, lonely, sad) 0.096 Moral (wrong, innocent, duty, decent, excuse) 0.102
$2^{nd}$ person pronouns (you, your, you’re, yourself, you’ve) 0.080 Female ( her, she, girl, she’s, mom) 0.060
Past focus (was, did, were, been, didn’t ) 0.069 Family (son, married, uncle, dad, mom) 0.055
All or none (no, all, nothing, never, yes) 0.055 Sexual (chaste, lust, sex, sexy, pimp) 0.051
Family (son, dad, baby, mom, mama) 0.054 Conflict (kill, killed, accusing, killing, cruel) 0.024
Social References (you, your, we, he, you’re) 0.050 Swear words (hell, bloody, idiot, damn, ass) 0.024
Anxiety (fear, afraid, worry, terrified, scared) 0.042 Anger (mad, angry, hate, cruel, argue) 0.022
Illness (sick, pain, pains, flu, sickly) 0.038 Present focus (is, are, don’t, i’m, aren’t) 0.021
Table 3: Pearson r for LIWC categories significantly correlated ($p<0.05$) with shame for Hollywood and Bollywood after Benjamini-Hochberg p-value correction (See Appendix for confidence intervals and p values). Note, $1^{st}$ person sing pronouns are strongly correlated with Hollywood-shame whereas the correlation with $2^{nd}$ person pronouns and social references are up to 3 times stronger in Bollywood compared to Hollywood. See Table A7 and A9 for the complete set of correlations.
3 Approach
Fig. 1 illustrates our study design comprising a blend of two key approaches (a) a psychosocial vocabulary approach to measure how expressions of shame and pride differ and (b) prompting LLM approach to extract reasons behind them and how they differ between India and the U.S.A.
3.1 Vocabulary Approach: LIWC
Linguistic Inquiry of Word Count (LIWC) Boyd et al. (2022) is a corpus analysis tool widely used in psychology to identify psychological categories of words such as "self-focus" (e.g., I, my, mine) and "other-focus" (e.g., you, yours, theirs) which are of great significance for examining social behaviors such as self-regulation and conformity.
To understand cross-cultural linguistic variations in the manifestation of shame and pride, we computed the normalized distribution of psycho-social categories from the dialogues and examined their correlation with shame and pride compared to the control set. The search keywords used for building social emotions corpus (See Table A1) were removed from the LIWC dictionary to prevent overestimation of shame- (e.g., negative emotion) and pride-related categories (e.g., achievement).
3.2 Pretrained LLM: GPT-4 Chat
Identifying why shame/pride is expressed in discourse requires multicultural world knowledge. LLMs such as GPT-4 exhibit a superlative pragmatic understanding of the world around us and are increasingly used for extracting implicit meanings and beliefs Pan et al. (2023); Törnberg (2023). We used GPT-4 chat in a two-shot setting ( See Tables A2 and A3 for prompts) to identify:
  • who is $<$being shamed/feeling proud$>$ in the given movie discourse, and what is their gender? and,
  • What is the reason behind $<$the feeling of shame/pride$>$?
Since there are at least two characters in a discourse, the first question orients the LLM to focus on the person experiencing the social sanction or approval and then identify their gender. The output for the second question serves as the implicit social expectation that led to the expression of shame/pride in the culture. Asking for "reason" leading to the expression of social emotion encourages LLMs to retrieve expectations from the provided context, mitigating potential Anglocentric tendencies Havaldar et al. (2023b).
Thematic Analysis of Reasons Behind Shame/Pride-related Expressions
To capture overarching themes in social expectations that led to shame/pride-related expressions in Indian and American societies, we performed agglomerative clustering after embedding unique shame and pride-related reasons using SBERT embeddings Reimers and Gurevych (2019). Unlike movie dialogues, reasons extracted from GPT-4 chat are short phrases (See Table 6) and are devoid of culture-specific language style markers (i.e., Indian English vs American English) thus, no control for culture-specific language is needed.
Hollywood: LIWC Category (Top-5 words) $r$ Bollywood: LIWC Category (Top-5 words) $r$
Social References (you, your, we, he, you’re) 0.131 Drive (we, our, us, sir, work) 0.122
Family (son, dad, baby, mom, mama) 0.121 Achieve (work, better, win, best, try) 0.098
$1^{st}$ person sing. pronouns (i, i’m, me, my, i’ll) 0.095 Power (sir, own, respect, kill, power) 0.091
$2^{nd}$ person pronouns (you, your, you’re, you’ve, yourself) 0.091 Social References (you, your, he, we, our) 0.085
Male References (he, his, him, man, son) 0.084 Moral (wrong, duty, brave, arrogant, useless) 0.075
Drive (we, our, us, we’re, dad) 0.076 Reward (win, won, glory, success, successful) 0.071
Ethnicity (American, Irish, Chinese, German, Christian) 0.052 Political (nation, army, sultan, president, dynasty) 0.064
Reward (win, won, winner, successful, earned) 0.048 $1^{st}$ person pl. pronouns (we, our, us, we’ll, let’s) 0.060
Achieve (work, better, best, trying, try) 0.045 Family (son, papa, married, dad, uncle) 0.059
Power (sir, own, war, strong, mighty) 0.044 Ethnicity (Indian, Indians, British, Hindi, caste) 0.054
Table 4: Pearson r for LIWC categories significantly correlated ($p<0.05$) with pride for Hollywood and Bollywood after Benjamini-Hochberg p-value correction (See Appendix for confidence intervals and p values). Note the contrast, Achievement-related pride in Bollywood, Self-centered pride vs We-centered pride, Family in Hollywood vs Family in Bollywood See Tables A8 and A10 for all significantly correlated categories.
Bollywood Hollywood
shame pride shame pride
male 2690 1259 591 776
female 1215 326 246 236
Not clear 1416 652 319 719
total 5321 2237 1156 1731
Table 5: Reason and Gender distribution (with duplicates) for dialogues for which GPT-4 predicted male or female. The duplicate reasons are not removed as their frequency reflects their prevalence and is useful for estimating gender association.
4 Results
4.1 How is shame/pride expressed across cultures?
Shame is associated with negative emotions, power, and morality in both movie industries (See Table 3), however, significant cultural nuances exist in its manifestation. In Hollywood, shame is more I-focused with remorse (e.g., sadness, anxiety) and past-focused language, whereas in Bollywood, it is more You-focused with frequent social references and present-focused language. In Bollywood, psychosocial categories associated with shame reveal a larger social context (note the contrast in family i.e., nuclear family in Hollywood vs extended family in Bollywood) and social situations (i.e., anger, swear, conflict). Categories female and sexual are exclusively correlated with shame in Bollywood potentially indicating an honor system in collectivist societies Caffaro et al. (2014) whereas it is illness that is uniquely correlated with shame in Hollywood indicating social sanctions around incapability.
A similar dichotomy is observed for pride (See Table 4). Pride in Hollywood is more centered toward family (note, family in Bollywood is male-centric and comprises extended family (such as uncle)) and self-centered (e.g., I am proud of __) whereas it is achievement/power focused and "We"-centered in Bollywood movies (with othering (us vs they)- See Table A8). Male references are more strongly associated with pride than females in Hollywood whereas the female category has an insignificant ($p>0.05$) correlation with pride in Bollywood. Pride is exclusively positive (see POSEMO in Table A10) in Hollywood whereas in Bollywood, we suppose that it relates to honor more (e.g., protecting family’s pride, bringing pride to family) and is not always used in a positive context.
4.2 How do social expectations behind shame/pride vary across cultures?
Cross-Cultural Shame/Pride Dataset
For the Bollywood set, GPT-4 chat predicted reasons (also implicit social expectations) for 5321 (98.4%) shame-related dialogues out of 5409, and 2237 (74.6%) pride-related dialogues out of 2999. For the Hollywood set, GPT-4 chat predicted a reason for 1156 (94.6%) shame-related dialogues out of 1221 and 1731 (61.7%) pride-related dialogues out of 2805. Upon manual analysis, we found that pride is also used to express affection, specifically toward close family members such as sons - therefore, a lower number of reasons is associated with pride compared to shame. GPT-4 chat predicted a total of 10,445 social expectations (See Table 5 for distribution). Prompting "reasons" behind shame and pride allowed us to capture high specificity in cultural norms (See Tables 6 for shame-related expectations and A6 for pride-related expectations).
4.2.1 Manual Evaluation
Two volunteers manually verified the predicted gender for "the person experiencing shame/pride" and the predicted "reason" in a randomly sampled set of $100$ dialogues each from Hollywood and Bollywood. The annotator who labeled the Bollywood set is an Indian, aware of social roles and expectations in Indian society. Likewise, the annotator for the Hollywood set is an American with a nuanced understanding of the social expectations of the U.S.A.
Bollywood
eavesdropping on private conversation
expressing love for a man
Incestuous relationship
giving birth to a girl child
Hollywood
not living up to expectations
hiding/avoiding confrontation
not returning calls after intimacy
mistreatment of a woman
Table 6: A subset of reasons extracted from movie dialogues expressing shame. A total of 4604 unique reasons (Bollywood-3660, Hollywood-944) were extracted.
For Bollywood, 8% of predicted gender and 11% of predicted reasons were incorrect where it was 5% and 2% respectively for Hollywood (See Table A5 for more details). There were 20 cases where gender (15) and/or reason (9) were ambiguous whereas, for Hollywood, there were 11 (gender=10, reason=1) such samples.
4.2.2 Thematic Analysis
Twenty-four clusters for shame-related reasons and fifteen for pride-related reaosns were formed using agglomerative clustering ( See Tables A13 and A14 for clustering parameters). The clusters were manually assigned a theme (as depicted on the Y-axis of Fig. 2 and 3) after analyzing the ten closest samples based on cosine distance within the cluster (See Table A11 and A12 for examples in each cluster). We computed the relative association for each cluster with Bollywood and Hollywood using eq. 1 and performed the Barnard-Exact Test Barnard (1947) with the Yates Correction Yates (1934) to test if the possibility of observing norms related to pre-assigned themes is statistically different across movie industries.
$\vec{\Delta}=\forall_{t_{i}\in themes}\frac{\mathcal{D}_{bolly_{t_{i}}}}{{\mathcal{D}_{bolly}}}-\frac{\mathcal{D}_{holly_{t_{i}}}}{\mathcal{D}_{holly}}$ (1)
Here, $\mathcal{D}_{industry_{t_{i}}}$ represents dialogues to $industry\in\{holly,bolly\}$ and ${t_{i}}\in themes$ (as depicted on Y-axis in Fig 2 and 3).
Figure 2: Relative association ($\Delta$) of Bollywood and Hollywood to themes obtained from agglomerative clustering performed on shame-related norms. * indicates significant difference i.e., $p>0.05$. See Table A15 for statistics and p-value.
Shame-related social norms
Themes such as lack of accountability and poverty are more common in Hollywood whereas inappropriate sexual behavior and gender roles are more prevalent in Bollywood (See Fig. 2). Unsurprisingly, collectivist factors such as non-conformity in gender roles, disrespect, and deviation from family norms are strongly associated with shame in Bollywood whereas individualistic attributes such as poverty, causing harm and incompetence evoke shame in Hollywood.
Pride-related social norms
Duty, doing the "right" thing, and, self-identity are associated with pride in Hollywood whereas Ethnolinguistic identity, and son’s achievements are associated with pride in Bollywood (See Fig. 3).
4.2.3 How is shame/pride attributed across genders?
For the Bollywood set, GPT-4 predicted 1541 targets as female and 3949 as male. For the Hollywood set, GPT-4 predicted 482 targets as female and 1367 as male. Across all combinations (shame vs pride x Bollywood vs. Hollywood in Table 5), there are more male targets than females in line with the findings of Geena Davis Inclusion Quotient [^4] https://about.google/intl/ALL_us/main/gender-equality-films/.
We computed the gender-wise attribution to "shame" and "pride" in movie dialogues using eq. 2. A positive score indicates a higher association of gender groups with pride, whereas a negative score reflects a higher association with shame. A null score indicates no preference.
Figure 3: Relative association ($\Delta$) of Bollywood and Hollywood to themes obtained from agglomerative clustering performed on pride-related norms. * indicates significant difference i.e., $p>0.05$. See Table A16 for statistics and p-value.
Figure 4: Relative association ($\Delta$) of emotions pride and shame gender-wise in Hollywood and Bollywood dialogues. A higher positive score indicates a stronger association of gender with pride.
$\vec{\Delta}_{g}=\forall_{g\in\{male,female\}}\frac{\mathcal{D}_{pride_{g}}-\mathcal{D}_{shame_{g}}}{\mathcal{D}_{g}}$ (2)
Here, $\mathcal{D}_{pride_{g}}$ represents movie dialogues having pride-related words where the target of pride is $g\in\{male,female\}$, $\mathcal{D}_{shame_{g}}$ represents movie dialogues having shame-related word where the target of shame is $g\in\{male,female\}$ and $\mathcal{D}_{g}$ represents total movie dialogues grouped based on gender. As depicted in Fig. 4, Hollywood movies are pride-oriented, whereas Bollywood movies are shame-oriented. Females are attributed more shame, and the difference (male-female) in the expression of pride and shame is $0.16$ for Hollywood and $0.21$ for Bollywood.
Social expectations evoking shame are similar across movie industries (See Fig. 5 and 6). Sexuality (e.g., promiscuity, immodesty) is the dominant theme for women in both movie industries whereas males are shamed for incompetency in both industries. Women across movie industries express more pride in family roles. Men-pride in Bollywood is centered on justice, winning and bravery whereas it is duty, self-identity and winning in Hollywood (See Fig. A1 for gender differences in pride).
Figure 5: Gender differences in shame-related themes in Bollywood movies. * indicates significant difference i.e., $p>0.05$.
Figure 6: Gender differences in shame-related themes in Hollywood movies. * indicates significant difference i.e., $p>0.05$.
5 Discussion
We release a cross-cultural dataset of shame (6477 samples) and pride (3968 samples) related expressions with implicit reasons behind them. To the best of our knowledge, the only related corpus is the GoEmotions dataset Demszky et al. (2020) comprising 817 samples for embarrassment and 452 samples for pride. This corpus can serve the task of developing computational models for detecting shame/pride expressions and recognizing how appraisals of such emotions vary -- enabling culturally cognisant human-computer conversations for diverse users Kim et al. (2011) and cross-cultural intent translation [^5] DARPA Computational Cultural Understanding: https://www.darpa.mil/program/computational-cultural-understanding.
Shame is a highly undesirable self-focused emotion emphasizing incompetency and failures and is rarely used in the U.S. Cohen (2003); Boiger et al. (2013) whereas Eastern cultures use shame more frequently and have a desirable affect Lim (2016). Notably, we also observe infrequent use of shame in Hollywood movies and self-focus (reflecting internal shame) (See Table 2 and 3 for frequency distribution and psychosocial constructs). Unlike in Western nations, shame is interdependent in collectivist communities Wong and Tsai (2007). The empirical analyses reveal other-focused (also, the public nature of shame) in Bollywood movies (See Table A7). Moreover, the contrasting tenses coupled with varying emotions, i.e., past + sadness in Hollywood vs. present + anger in Bollywood, reflect their varying goals, i.e., remorse for past failures/losses vs enforcing conformity Wong and Tsai (2007). Pride-related discourse in Hollywood is duty and achievement-focused, in line with prior findings underlining the increasing significance of "success" with growing individualism Cohen (2003), whereas in Bollywood, pride is centered around collective achievement (we, achievement, prosocial). It is also interesting to note linguistic markers such as determiner (used with objects/nouns) in Bollywood indicating more materialistic pride, compared to personal pronouns and social references in Hollywood reflecting more personal pride (See Tables A8, A10 for all significant categories correlated with pride).
Eliciting reasons leading to shame and pride expressions revealed social expectations across cultures. We observed high specificity in extracted reasons revealing cultural subtleties (e.g., desire for a son in Bollywood vs. returning calls after date night in Hollywood). To tease apart ways in which cultures differ, we first considered mapping social expectations/generated reasons to Schwartz’s Theory of Values Schwartz (2012). However, social expectations were found to have contrasting values depending on the culture. Consider "refusing to marry", an instance of non-conformity in Indian society, whereas it is an instance of self-direction in a Western context. To effectively understand the cross-cultural variations in social expectations underlying shame/pride, we performed hierarchical clustering and empirically picked the distance after manually analyzing the quality and granularity of clusters. The situations connected to shame and pride in the U.S. society (i.e., lacking accountability, poverty, harm) vs Indian society (e.g., inappropriate sexual behavior, gender roles, betrayal - see Fig 2 and 3) align with known cultural tendencies (i.e., individualist vs collectivist) of both nations Triandis (1988).
The vision of safe and accountable AI is centered on LLMs’ moral and value alignment. The dominant approach for norm discovery involves prompting LLMs, sometimes coupled with a verification step such as an entailment test or underlying emotion (negative emotion $\rightarrow$ norm violation) Jiang et al. (2021); Fung et al. (2022); CH-Wang et al. (2023). However, most of the social situations and the human annotators employed to label those situations reflect English beliefs and ethics. Our approach to extracting culture-specific social expectations and attitudes using social emotions overcomes this limitation. While the clustering was performed on the reasons extracted from movie conversations, the cluster themes are similar to patterns seen during LIWC analysis, validating our approach for overcoming Anglocentric bias in LLMs for norm discovery Ziems et al. (2022).
Lastly, we examined how social expectations associated with shame/pride are attributed to men and women across cultures influencing their social behaviors; women are less assertive across cultures to avoid negative attribution Amanatullah and Morris (2010); Ferguson and Eyre (2000) and ask for less during negotiation than men Arnold and McAuliffe (2021). We demonstrated that shame, a negative self-conscious emotion expressing devaluation, is targeted towards women more than men, whereas pride, endorsing increased social value, is used for more men than women in movies. The similarity in social expectations (e.g. regulating sexual behaviors in women vs shaming incompetency in men) across cultures is surprising (See Fig 5 and 6). It is important to characterize the source and acquired social expectations before their use for achieving alignment in LLMs otherwise we risk social bias.
6 Conclusion
We introduce a multi-cultural dataset of shame/pride-related expressions and the underlying social expectations. This study demonstrates (a) cross-cultural linguistic differences in shame and pride-related expression, offering insights into their functions across cultures, (b) cultural dichotomy in social expectations, and (c) more "social sanctions" and fewer "endorsements" to the female gender in movies. Future work can utilize our dataset for culturally aligned LLMs and build social emotion perception and appraisal in human-computer interactions. Additionally, our analysis of shame/pride is the first of its kind; we hope future NLP researchers will build upon this work to investigate social expectations in LLMs from a multicultural, social emotion-based perspective.
Social Impact and Ethics Statement
Social norms discovery is a crucial component in programs [^6] ALIGN-https://www.alignplatform.org/learning-collaborative designed for social and behavioral change to promote equity, social justice, and well-being Mauduy et al. (2022); Bonan et al. (2020). Further work explores style as a product of norm differences Havaldar et al. (2023a). Social psychology investigates social norms (descriptive vs injunctive) to design experiments for understanding behaviors such as self-regulation, persuasion Cialdini et al. (1990) and decision-making Gavrilets (2020); Bhanot (2021) to promote collective-level change in societies.
Relatedly, Kimbrough and Vostroknutov (2023) showed people’s tendency to choose self-serving social norms using a dictator-recipient setup, emphasizing the need for dedicated research efforts to understand morality and belief distortion in different contexts. The norms and cultural preferences learned from movies that often showcase stereotypical behaviors of society may induce pluralistic ignorance and, more importantly, lead to discrimination and biases in LLMs when used for training. We hope that this paper will encourage scrutiny of source corpora and derived norms before their use for fine-tuning LLMs.
Limitations
Social norms mutate as society evolves. We acknowledge that our dataset of movies (released post-1990) may reflect social norms that are less characteristic of contemporary society. Moreover, countries like India and America contain a mix of cultures. The captured norms may not reflect the cultural variations, for example, between regions (e.g., East Coast vs West Coast in the U.S.A or North India vs South India). Movies also exaggeratedly depict the world around us (e.g., weddings, criminal activities, sexual abuse, etc.), and we caution against stereotyping cultures based on movie-based norms.
The dominant language in Bollywood movies is Hindi and our analysis is based on their English translations which may not always be accurate, especially when the discourse is about concepts native to a culture. Relatedly, the LIWC may not have high coverage for such concepts. We did not compare the movie genre and acknowledge that situational/unrealistic norms (e.g., science fiction, comedy, etc.) could exist. Regardless, social norms associated with shame and pride are still relatable and reflect the target audience’s beliefs. It is also worth noting that gender norms in movies may not fully reflect the norms of today’s society. This study is conducted over a thirty-year long period and thus, the aggregated norms may not reflect the current trends.
We acknowledge that the excerpts expressing shame and pride from movie subtitles may have led to the over-representation of certain social situations (e.g., son’s achievement and daughter’s wedding in India vs duty and competence in America). The prompts used to elicit norms may have induced unwanted bias Cheng et al. (2023); Lucy and Bamman (2021) and it is worth investigating the variations if any, in extracted norms due to different prompt designs.
Data use and availability
Bollywood subtitles were scraped from a website called www.bollynook.com and Hollywood subtitles are taken from publicly available Kaggle movie subtitles datasets. We will only release dialogue excerpts related to shame and pride. We have around 7.5k such instances in Bollywood and almost 2.8k in Hollywood. Under the interpretation of fair use in research, this makes up a very small portion of dialogues taken from Bollywood (i.e. almost 3 samples (5 lines long) per movie on average ) and Hollywood (about 1 sample (5 lines long) per movie on average).
Acknowledgements
We acknowledge the use of ChatGPT, an AI language model developed by OpenAI, for its assistance in refining this manuscript.
Appendix A Appendix
A.1 Search Keywords
Search keywords defined in Table A1 were used to extract shame- and pride-related dialogues from Bollywood and Hollywood movie subtitles.
Lemma Word forms
shame shame, shamed, shameful, ashamed
pride proud, prouder, proudly, pride
Table A1: List of search keywords for extracting dialogues.
A.2 Examining Implicit shame
We prompted GPT-4 chat to identify (a) if a given discourse expresses shame, and (b) identify the behavior or action behind shame, for a randomly sampled set of 100 samples without explicit mentions of shame/pride each from Hollywood and Bollywood. We asked two human annotators to evaluate the correctness of GPT-4 chat-generated outputs.
A.3 Prompts
The prompts designed to extract the subject experiencing self-conscious emotions, their gender, and the reason behind the emotions are provided in Tables A2 and A3.
Prompts for norm discovery in Bollywood movie dialogues
"You are an experienced social scientist and you study Indian society. Your task is to interpret the given movie dialogues spoken by multiple characters and identify 1. Who is feeling proud? Provide the gender. If agent or gender is not explicit, use the pronouns, name, and salutations to guess agent and gender. 2. Identify the reason behind feeling proud. If not explicit, return NA. The answer should be short and in below CSV format. $<$who_proud, gender, reason$>$
Input: Your mother had already given up on me because I was useless hopeless But you were the apple of her eye My child will make me proud look after me He’ll support in my old-age But this kid left her all alone.
Output: mother, male, provide care for old parents Input: I’ve heard there’s a promising young student in your school What’s his name? He’s made us proud in long jump, we are here to felicitate him Call him Show yourself, Raju Tempre Output: authority, NA, Sports achievement"
"You are an experienced social scientist and you study Indian society. Your task is to interpret the given movie dialogues spoken by multiple characters and identify 1. Who is being shamed? Predict the gender. If gender is not explicit, use the pronouns, name, and salutations to guess gender. 2. Identify the primary reason for shaming The answer should be short and in CSV format. $<$ shamed, gender, reason$>$ Input: And should we bow before others begging....them to marry our daughters? This shall not happen. Neither will the girls be alive here nor shall....we be ashamed of ourselves. You cannot kill the life which God has given. I won’t let you commit the sin. Output: girl’s parent, NA, not able to marry off their daughters Input: Black marketers are now in the open. And the thieves too Politics is in a great mess Shame on this system. There’s no democracy Get rid of these politicians The gong has struck..’’Our hearts are swaying to it’s beats’’ Output: System, NA, poor law and regulations"
Table A2: Bollywood: Prompts for norm discovery using GPT-4 Chat. The temperature was set to 00 to minimize randomness.
Prompts for norm discovery in Hollywood movie dialogues
"You are an experienced social scientist and you study American society. Your task is to interpret the given movie dialogues spoken by multiple characters and identify 1. Who is feeling proud? Provide the gender. If agent or gender is not explicit, use the pronouns, name, and salutations to guess agent and gender. 2. Identify the reason behind feeling proud? If not explicit, return NA. The answer should be short and in below CSV format. $<$who_proud, gender, reason$>$ Input: I want to go to Worlds and win gold. I want to go to the 88 Olympics in Seoul and win gold. Good! I’m proud of you. Are you getting the support that you need? What do you mean sir? Output: Sir, male, winning olympic gold Input: Yes. Yes, I did. I promise, this time I really got the promotion. - I’m proud of you, son. - Thank you, sir. Excuse me. Hi, sweetheart. Output: father, male, for getting the promotion"
"You are an experienced social scientist and you study American societies. Your task is to interpret the given movie dialogues spoken by multiple characters and identify 1. Who is being shamed? Predict the gender. If gender is not explicit, use the pronouns, name, and salutations to guess gender. 2. Identify the primary reason for shaming. The answer should be short and in CSV format. $<$ shamed, gender, reason, prevalence $>$ Input: You still owe me 100. Remember? You stiffed Donny for 100 bucks? Cheapskate. Shame on you. Pay this man his C-note. Now I know why they call you the Snake. Output: NA, male, not returning borrowed money Input: You prey on your own people. You steal from your own people. Have you no shame!? - Huh? - Well, we’re still here. Man: Mr. Markopolos, it’s all yours. Output: Snake, male, stealing and preying on people"
Table A3: Hollywood: Prompts for norm discovery using GPT-4 Chat. The temperature was set to 00 to minimize randomness.
A.3.1 Annotation
The annotation guidelines to verify the gender predicted by GPT-4 and the correctness of the reason are provided in Table A4. The annotators for Bollywood set and Hollywood set were Indian and American respectively. Both annotators were female, proficient in English language, and well-versed with social norms. During annotation, if the gender or the reason is unclear, the annotators were asked to label "not explicit". We only considered the cases where the gender was predicted to be either male or female. The task is objective and inter-annotator agreement was not computed. The annotators volunteered for the task and were not provided monetary compensation.
Guidelines for Manual Evaluation
Step-1. Read the conversation and identify the person feeling ashamed (or being shamed) or proud.
Step-2. Identify the gender. Check gender markers such as Mr/Mrs., s/he, him/her, etc. If the name is provided in the conversation, check if the name is likely to be a male name or female. If not clear, mark "not explicit".
Step-3. Read the reason behind shame/pride. Compare with conversation and determine if the provided reason is the cause for shame/pride.
Table A4: Guidelines for Annotation
Gender Evaluation
Incorrect Ambiguous
Bollywood 8 15 (3 Female)
Hollywood 5 10 (3 Female)
Social Norms/Reason Evaluation
Incorrect Ambiguous
Bollywood 11 9
Hollywood 2 1
Table A5: Manual Evaluation of predicted gender and reasons for randomly sampled 100 samples each from Bollywood and Hollywood.
Bollywood
mastering a trick
fulfilling father’s dreams
provide care for old parents
fiancee’s physical appearance
his wealth
Hollywood
for being a hard worker regardless of the task
being a brilliant student
winning Olympic gold
achievements and growth
coming out as queer
Table A6: A subset of reasons extracted from movie dialogues expressing pride. A total of 3163 unique reasons (Bollywood-1589, Hollywood-1574) were extracted.
A.4 LIWC Correlation Results
Tables A7, A8, A9 and A10 contain the significantly correlated ($p<0.05$) LIWC categories, the most frequent five words for each category, pearson r and 95% confidence interval.
LIWC Categories Top-5 words r p-value 95% CI
Negative emotions (EMO_NEG) (bad, mad, scared, worry, fear) 0.330 0.000 [0.315, 0.344]
Negative tone (TONE_NEG) (lost, kill, wrong, bad, hit) 0.249 0.000 [0.233, 0.265]
POWER (sir, respect, own, kill, poor) 0.198 0.000 [0.182, 0.214]
EMOTION (love, good, bad, happy, crazy) 0.168 0.000 [0.152, 0.185]
YOU (you, your, you’re, yourself, you’ve) 0.161 0.000 [0.145, 0.178]
Social references (SOCREFS) (you, your, he, her, him) 0.148 0.000 [0.132, 0.165]
FEELING (feel, touch, feeling, felt, hard) 0.133 0.000 [0.117, 0.150]
DRIVES (we, our, us, sir, married) 0.115 0.000 [0.098, 0.132]
SOCIAL (you, your, he, her, him) 0.110 0.000 [0.094, 0.127]
MORAL (wrong, innocent, duty, decent, excuse) 0.102 0.000 [0.085, 0.118]
AFFECT (love, good, keep, respect, well) 0.078 0.000 [0.062, 0.095]
NEGATE (not, don’t, no, aren’t, won’t) 0.065 0.000 [0.049, 0.082]
FEMALE (her, she, girl, she’s, mom) 0.060 0.000 [0.043, 0.077]
Personal Pronouns (PPRON) (you, i, me, your, my) 0.058 0.000 [0.041, 0.075]
FAMILY (son, married, uncle, dad, mom) 0.055 0.000 [0.039, 0.072]
Preposition (PREP) (to, of, in, for, on) 0.053 0.000 [0.037, 0.070]
SEXUAL (chaste, lust, sex, sexy, pimp) 0.051 0.000 [0.034, 0.067]
PRONOUN (you, i, me, your, my) 0.040 0.000 [0.023, 0.057]
AUXVERB (is, are, have, be, don’t) 0.029 0.001 [0.013, 0.046]
CONFLICT (kill, killed, accusing, killing, cruel) 0.024 0.007 [0.007, 0.041]
SWEAR (hell, bloody, idiot, damn, ass) 0.024 0.007 [0.007, 0.041]
Anger (EMO_ANGER) (mad, angry, hate, cruel, argue) 0.022 0.015 [0.005, 0.039]
FOCUSPRESENT (is, are, don’t, i’m, aren’t) 0.021 0.018 [0.005, 0.038]
Table A7: Psychosocial categories significantly correlated (p<0.05) with shame in Bollywood dialogues. p-values were corrected using Benjamini-Hochberg correction. The categories are arranged in decreasing order of correlation.
LIWC Categories Top-5 words r p-value 95% CI
DRIVES (we, our, us, sir, work) 0.122 0.000 [0.104,0.140]
Determiner (DET) (the, a, my, your, that) 0.101 0.000 [0.082,0.119]
ACHIEVE (work, better, win, best, try) 0.098 0.000 [0.080,0.116]
POWER (sir, own, respect, kill, power) 0.091 0.000 [0.072,0.109]
Social References (SOCREFS) (you, your, he, we, our) 0.085 0.000 [0.067,0.103]
Preposition (PREP) (of, to, in, for, with) 0.084 0.000 [0.066,0.103]
MORAL (wrong, duty, brave, arrogant, useless) 0.075 0.000 [0.057,0.093]
Conjunction (CONJ) (and, but, so, if, as) 0.075 0.000 [0.056,0.093]
REWARD (win, won, glory, success, successful) 0.071 0.000 [0.052,0.089]
Positive tone (TONE_POS) (love, good, thank, well, great) 0.064 0.000 [0.046,0.082]
POLITIC (nation, army, sultan, president, dynasty) 0.064 0.000 [0.045,0.082]
WE (we, our, us, we’ll, let’s) 0.060 0.000 [0.042,0.079]
SOCIAL (you, your, he, we, our) 0.060 0.000 [0.041,0.078]
FAMILY (son, papa, married, dad, uncle) 0.059 0.000 [0.041,0.078]
AFFILIATION (we, our, us, dear, we’ll) 0.059 0.000 [0.041,0.077]
FEELING (feel, feeling, hard, felt, sense) 0.055 0.000 [0.036,0.073]
ETHNICITY (indian, indians, british, hindi, caste) 0.054 0.000 [0.036,0.072]
MALE (he, his, him, son, sir) 0.054 0.000 [0.036,0.072]
CULTURE (indian, nation, army, car, indians) 0.044 0.000 [0.026,0.062]
AFFECT (love, good, thank, well, great) 0.040 0.000 [0.021,0.058]
ARTICLE (the, a, an, that) 0.039 0.000 [0.021,0.058]
Personal Pronouns (PPRON) (you, i, my, your, me) 0.036 0.000 [0.017,0.054]
PROSOCIAL (thank, please, sorry, respect, gift) 0.032 0.001 [0.014,0.051]
FUNCTION (you, the, i, of, to) 0.032 0.001 [0.013,0.050]
YOU (you, your, you’re, you’ve, you’ll) 0.030 0.003 [0.011,0.048]
THEY (they, their, them, they’re, they’ll) 0.029 0.003 [0.011,0.048]
CERTITUDE (really, real, surely, proved, actually) 0.021 0.041 [0.002,0.039]
Positive Emotion (EMO_POS) (love, good, happy, happiness, smile) 0.020 0.049 [0.001,0.038]
Table A8: Psychosocial categories significantly correlated (p<0.05) with pride in Bollywood dialogues. p-values were corrected using Benjamini-Hochberg correction. The categories are arranged in decreasing order of correlation.
LIWC Categories Top-5 words r p 95% CI
Negative emotions (EMO_NEG) (sick, pain, fear, bad, afraid) 0.425 0.000 [0.403, 0.446]
Negative tone (TONE_NEG) (lost, wrong, sick, pain, poor) 0.355 0.000 [0.331, 0.377]
EMOTION (good, love, sick, pain, bad) 0.290 0.000 [0.266, 0.314]
POWER (own, sir, poor, killed, war) 0.263 0.000 [0.238, 0.287]
AFFECT (well, good, love, help, damn) 0.168 0.000 [0.142, 0.193]
DRIVES (we, us, our, work, we’re) 0.152 0.000 [0.127, 0.178]
FUNCTION (you, i, the, to, of) 0.131 0.000 [0.105, 0.157]
Personal Pronouns (PPRON) (you, i, me, i’m, my) 0.111 0.000 [0.085, 0.137]
MORAL (wrong, excuse, decent, honest, duty) 0.110 0.000 [0.084, 0.136]
I (i, me, i’m, my, i’ll) 0.102 0.000 [0.076, 0.128]
Sadness (EMO_SAD) (crying, cry, sob, lonely, sad) 0.096 0.000 [0.070, 0.122]
NEGATE (no, not, don’t, nothing, never) 0.090 0.000 [0.063, 0.116]
Preposition (PREP) (to, of, in, for, on) 0.089 0.000 [0.062, 0.115]
PRONOUN (you, i, that, it, me) 0.088 0.000 [0.062, 0.114]
YOU (you, your, you’re, yourself, you’ve) 0.080 0.000 [0.054, 0.106]
Auxiliary verb (AUXVERB) (be, i’m, is, was, have) 0.077 0.000 [0.051, 0.103]
FOCUSPAST (was, did, were, been, didn’t) 0.069 0.000 [0.043, 0.096]
SOCIAL (you, your, we, he, you’re) 0.060 0.000 [0.033, 0.086]
Conjunction (CONJ) (and, so, but, if, when) 0.056 0.000 [0.030, 0.082]
LINGUISTIC (you, i, the, to, of) 0.056 0.000 [0.030, 0.082]
ALLNONE (no, all, nothing, never, yes) 0.055 0.000 [0.029, 0.081]
FAMILY (son, dad, baby, mom, mama) 0.054 0.000 [0.028, 0.081]
Social references (SOCREFS) (you, your, we, he, you’re) 0.050 0.000 [0.024, 0.077]
Anxiety (EMO_ANX) (fear, afraid, worry, terrified, scared) 0.042 0.004 [0.015, 0.068]
ILLNESS (sick, pain, pains, flu, sickly) 0.038 0.010 [0.012, 0.064]
FEELING (feel, felt, pain, feeling, hard) 0.035 0.020 [0.008, 0.061]
Anger (EMO_ANGER) (hate, hated, mad, angry, hates) 0.033 0.026 [0.007, 0.059]
DIFFER (not, but, if, didn’t, or) 0.033 0.027 [0.007, 0.059]
Discrepancy (DISCREP) (should, can, would, can’t, want) 0.032 0.029 [0.006, 0.059]
COGNITION (no, not, all, know, but) 0.032 0.033 [0.005, 0.058]
Table A9: Psychosocial categories significantly correlated (p<0.05) with shame in Hollywood dialogues. p-values were corrected using Benjamini-Hochberg correction. The categories are arranged in decreasing order of correlation.
LIWC Categories Top-5 words r p-value 95% CI
Personal Pronouns (PPRON) (you, i, i’m, me, my) 0.135 0.000 [0.113,0.158]
Social References (SOCREFS) (you, your, we, he, you’re) 0.131 0.000 [0.108,0.154]
FAMILY (son, dad, baby, mom, mama) 0.121 0.000 [0.099,0.144]
Conjunction (CONJ) (and, so, but, as, if) 0.114 0.000 [0.091,0.137]
SOCIAL (you, your, we, he, you’re) 0.102 0.000 [0.079,0.125]
FUNCTION (you, i, the, of, to) 0.100 0.000 [0.077,0.123]
I (i, i’m, me, my, i’ll) 0.095 0.000 [0.072,0.118]
YOU (you, your, you’re, you’ve, yourself) 0.091 0.000 [0.068,0.114]
MALE (he, his, him, man, son) 0.084 0.000 [0.061,0.107]
DRIVES (we, our, us, we’re, dad) 0.076 0.000 [0.053,0.099]
Positive tone (TONE_POS) (good, well, thank, great, love) 0.072 0.000 [0.049,0.095]
PRONOUN (you, i, i’m, that, it) 0.072 0.000 [0.048,0.095]
Auxiliary Verb (AUXVERB) (i’m, be, is, was, have) 0.071 0.000 [0.048,0.094]
Positive Emotion (EMO_POS) (good, love, happy, hope, wonderful) 0.070 0.000 [0.047,0.093]
Preposition (PREP) (of, to, in, for, on) 0.070 0.000 [0.047,0.093]
AFFILIATION (we, our, us, we’re, dad) 0.054 0.000 [0.031,0.078]
EMOTION (good, love, happy, hope, bad) 0.052 0.000 [0.029,0.075]
ETHNICITY (american, irish, chinese, german, christian) 0.052 0.000 [0.029,0.075]
REWARD (win, won, winner, successful, earned) 0.048 0.000 [0.025,0.071]
ACHIEVE (work, better, best, trying, try) 0.045 0.000 [0.022,0.068]
POWER (sir, own, war, strong, mighty) 0.044 0.001 [0.021,0.067]
AFFECT (good, well, thank, great, love) 0.041 0.001 [0.018,0.064]
MORAL (wrong, excuse, hero, brave, dignity) 0.040 0.002 [0.017,0.064]
FEMALE (her, she, she’s, girl, ladies) 0.034 0.009 [0.011,0.057]
CULTURE (american, car, president, nation, mayor) 0.033 0.012 [0.009,0.056]
FOCUSPAST (was, did, been, were, had) 0.032 0.013 [0.009,0.055]
SHEHE (he, his, him, her, she) 0.031 0.016 [0.008,0.054]
WORK (work, job, school, deal, company) 0.028 0.033 [0.005,0.051]
Table A10: Psychosocial categories significantly correlated (p<0.05) with pride in Hollywood dialogues with 95% confidence intervals. p-values were corrected using Benjamini-Hochberg correction. The categories are arranged in decreasing order of correlation.
A.5 Clustering Results
Tables A13 and A14 contain the manually annotated Cluster Theme, the total number of samples in each cluster, and Bollywood vs Hollywood distribution. The distance was set to $5$ and the duplicates were removed. The theme and top three examples demonstrating its meaning are provided in Tables A11 and A12.
Theme Examples from cluster
Stealing shamelessness and taking money, lying and misusing money, Stealing and bribery
Poverty not having money, being poor, living in poor conditions
Incompetence not fulfilling responsibilities, not meeting expectations or making a mistake, Failure or perceived incompetence
Promiscuity being promiscuous, Having a love affair, shameless behavior and expressing love
Inappropriate social behavior not living up to someone’s expectations, being disloyal, Being rude and ungrateful
Disobedience not upholding cultural values, disobedience or lack of respect, being disloyal
Parent-related not standing up for his mother, not taking care of his father, not taking responsibility for his son
Marriage-related refusing to marry, being forced into marriage, not accepting the proposed marriage
Immodesty Wearing inappropriate clothes, inappropriate behavior nudity in public, behaving indecently in public
Gender Roles not behaving as per her husband’s expectations, behaving inappropriately in front of her daughter, going against her husband
Sexual Harassment assaulting a girl, physical assault on women, sexual assault
Cowardice lack of courage, cowardice and inability to stand up for oneself, lack of pride and integrity
Alcoholism excessive drinking, inappropriate behavior due to alcohol, drinking and irresponsible behavior
Non-conformity not conforming to gender norms, breaking gender norms, behaving inappropriately according to societal norms
Illegal activities engaging in criminal activities, Being a criminal, committing illegal acts
Betrayal Betrayal or dishonesty, betrayal and infidelity, Deception/Betrayal
Lack of accountability not taking responsibility and blaming others, wrongdoing without remorse, not acknowledging wrongdoing
Disrespect disrespecting others, disrespecting an authority figure, Disrespectful and inappropriate behavior
Harm causing harm to others, committing harmful deeds, causing trouble and endangering others
Lying/Deception lying and hiding information, Lying or deceit, Deception/Not being truthful
Inappropriate sexual behavior inappropriate advances and comments, inappropriate behavior towards a young girl, inappropriate language and behavior
Social Etiquette Being humiliated in public, being disrespected and belittled, being mocked and treated shamefully
Privacy-related invading personal space, invading someone’s privacy, intrusion of personal space, trespassing/invading personal space
Accusation accused of wrongdoing, being accused of something, being accused of infidelity
Family norms Disrespecting family, causing harm and shame to family, causing difficulty and shame to family
Lack of shame perceived shamelessness, being perceived as shameless, feeling embarrassed and ashamed
Table A11: Themes ($n=26$) and top-3 examples for clusters obtained after agglomerative clustering of shame-related dialogues from Bollywood and Hollywood. The cluster Lack of shame was removed since the reason behind shame was not evident. Clusters Lack of accountability and Inappropriate social behavior were merged together (and named ‘‘Lack of accountability’’) due to overlapping normative expectations.
Theme Examples from cluster
Achievement being a brilliant student, working hard and achieving a spectacular result, professional achievement
Family Roles raising their child, having an unselfish mother, having a simple and good daughter
Self-identity being a man, being a woman, being gay
Duty fulfilling duty, Accomplishing something great, saving the world
Bravery bravery and sacrifice, bravery and service to nation, bravery and selflessness
Nation saving the country’s pride, making country proud, Making country proud
Doing the "right" thing for doing the right thing, for being heroes, for having courage
Son’s achievements son’s success and progress, son’s hard work and achievement, son’s determination
Family Honor maintaining dignity and pride, taking care of family pride, saving family’s pride
Justice standing up for justice, fighting for justice, bringing justice
Wedding marrying with pride and respect, sister’s marriage, daughter’s marriage into a reputed family
Physical Appearance & Assets his manhood, his wealth and beautiful wife, his knowledge and power
Resilience resilience and determination, overcoming struggles, enduring hardships without complaint
Winning Sports achievement, working hard and achieving a spectacular result, winning something
Ethnolinguistic Identity being Indian, defending the pride of Rajputs, being a Maharashtrian
Table A12: Themes ($n=15$) and top-3 examples for clusters obtained after agglomerative clustering of pride-related dialogues from Bollywood and Hollywood.
Theme (Shame-related norms) Total Samples Hollywood Bollywood
Lack of Accountability 802 185 617
Accusation 70 0 70
Alcoholism 98 18 80
Betrayal 262 26 236
Cowardice 209 51 158
Disobedience 216 49 167
Disrespect 246 26 220
Family norms 203 24 179
Gender roles 221 17 204
Harm 225 64 161
Illegal activities 253 27 226
Immodesty 311 58 253
Incompetence 342 83 259
Lying/Deception 222 53 169
Marriage-related 179 25 154
Non-conformity 208 28 180
Parent-related 263 34 229
Poverty 414 108 306
Privacy-related 102 14 88
Promiscuity 595 89 506
Inappropriate Sexual behavior 407 33 374
Sexual Harassment 232 24 208
Social Etiquette 417 97 320
Stealing 351 49 302
Total 6848 1182 5666
Table A13: Distribution of reasons (shame) across manually labeled clusters. A total of twenty-six clusters were generated with distance=5. Duplicates were removed for clustering. One cluster with generic reasons ( such as phrases "lack of shame") was removed and two clusters with similar reasons (related to accountability) were merged. Finally, 24 clusters were considered. The total is slightly more than numbers in Table 5 since a reason could be mapped to multiple clusters.
Theme (Pride-related norms) Total Samples Hollywood Bollywood
Achievement 332 174 158
Bravery 248 69 179
Doing the "right" thing 82 82 0
Duty 792 449 343
Ethnolinguistic Identity 148 7 141
Family Honor 263 77 186
Family Roles 356 156 200
Justice 255 109 146
Nation 216 57 159
Physical Appearance 106 49 57
Resilience 180 68 112
Self-identity 273 157 116
Son’s Achievements 315 87 228
Wedding 120 23 97
Winning 411 205 206
Total 4097 1769 2328
Table A14: Distribution of reasons (pride) across manually labeled clusters. A total of fifteen clusters were generated with distance=5. The total is slightly more than the numbers in Table 5 since a reason could be mapped to multiple clusters
(a) Pride-related norms in Bollywood
(b) Pride-related norms in Hollywood
Figure A1: Gender-wise differences in themes associated with normative expectations from pride-related discourse in Bollywood and Hollywood. $\Delta(Female-Male)$ is the difference between normalized dialogues attributed to females under a theme and to males under the same theme (as in eq. 1 but for dimension gender). * indicates significant difference i.e., $p>0.05$
Theme Statistic p-value
Inappropriate Sexual behavior 5.0380 0.00001
Gender roles 3.8262 0.0006
Betrayal 3.2044 0.001
Illegal activities 2.8258 0.005
Disrespect 2.8284 0.005
Sexual Harassment 2.8358 0.005
Promiscuity 1.5553 0.1
Accusation 3.8411 0.0006
Stealing 1.6798 0.09
Parent-related 1.8961 0.06
Family norms 2.0813 0.04
Non-conformity 1.4724 0.1
Marriage-related 1.1817 0.2
Privacy-related 0.9518 0.3
Alcoholism -0.2920 0.8
Immodesty -0.6634 0.5
Disobedience -2.1438 0.03
Lying/Deception -2.6508 0.008
Cowardice -2.7746 0.006
Incompetence -3.5187 0.002
Social Etiquette -3.3461 0.002
Harm -4.5140 0.00009
Poverty -4.9028 0.00002
Lack of accountability -4.6312 0.00007
Table A15: Barnard Exact test with Yates correction for testing statistically significant occurrence of shame-related themes in Bollywood and Hollywood.
Theme Statistic p-value
Ethnolinguistic Identity 9.6185 0.0000
Son’s Achievements 5.8025 0.0000
Bravery 5.0368 0.0000
Family Honor 4.7046 0.0000
Nation 5.1183 0.0000
Wedding 5.3898 0.0000
Resilience 1.4959 0.1370
Justice 0.1441 0.8963
Family Roles -0.2560 0.8024
Physical Appearance -0.6420 0.5223
Winning -2.8913 0.0038
Achievement -3.5425 0.0004
Self-identity -4.9482 0.0000
Doing the "right" thing -10.4936 0.0000
Duty -8.5488 0.0000
Table A16: Barnard Exact test with Yates correction for testing statistically significant occurrence of pride-related themes in Bollywood and Hollywood.